### v1

Abstract
Triangle counting in a graph is a fundamental problem and has a wide range of applications in various domains. It is crucial in understanding the structural properties of a graph and is often used as a building block for more complex graph analytics. In this paper, we solve the triangle counting problem in an anonymous graph in a distributed setting using mobile agents and subsequently use this as a subroutine to tackle the truss decomposition and triangle centrality problem. The paper employs mobile agents, placed on the nodes of the graph to coordinate among themselves to solve the triangle enumeration problem for the graph. Following the literature, we consider the synchronous systems where each robot executes its tasks concurrently with all others and hence time complexity can be measured as the number of rounds needed to complete the task. The graph is anonymous, i.e., without any node labels or IDs, but the agents are autonomous with distinct IDs and have limited memory. Agents can only communicate with other agents locally i.e., if and only if they are at the same node. The goal is to devise algorithms that minimise both the time required for triangle counting and the memory usage at each agent. We further demonstrate how the triangle count obtained through the mobile agent approach can be leveraged to address the truss decomposition, triangle centrality and local clustering coefficient problems, which involves finding maximal sub-graphs with strong interconnections. Truss decomposition helps in identifying maximal, highly interconnected sub-graphs, or trusses, within a network, thus, revealing the structural cohesion and tight-knit communities in complex graphs, facilitating the analysis of relationships and information flow in various fields, such as social networks, biology, and recommendation systems.
Keywords:
Mobile Agents Triangle Counting $k$-Truss Truss Decomposition Triangle Centrality Local Clustering Coefficient Time Complexity Memory Complexity Network Algorithms Distributed Algorithms
1 Introduction
Counting and listing triangles in a graph has received much attention in the last couple of decades as it serves as a building block of complex network analysis [54, 7]. The number of triangles in a graph is used for computing the clustering coefficient, one of the most used metrics for network analysis [54, 7], and triangle centrality [37, 1, 8]. Triangle counting also plays a pivotal role in the hierarchical decomposition of a graph such as truss decomposition [53] which is an important hierarchical subgraph structure in community detection [24, 2]. Triangle counting is used in solving many practical applications. Becchetti et al.  [6] used triangle counts in detecting web spam and estimating the content quality of a web page. Eckmann and Moses [18] have used the clustering coefficient in finding common topics on web pages. Other applications of triangle counting include query optimization in databases [5], link prediction in social network [51], and community detection in system biology [27].
In this paper, we are interested in the triangle counting problem along with its applications in truss decomposition, computing triangle centrality and local clustering coefficient using autonomous agents (or robots) on anonymous graphs. Suppose we are given $n$ agents positioned initially on the nodes (each node has one agent) of an $n$-node anonymous graph $G$. The agents coordinate among themselves to solve the triangle counting problem such that each agent (at a node) outputs (i) node-based triangle count, (ii) edge-based triangle count and the total number of triangles in the graph (see the problem statements in Section 3).
Our agent-based model has been gaining significant attention recently. For example, there have been some recent works on how to position the agents on nodes of the graph $G$ such that each agent’s position collectively form the maximal independent set (MIS) of $G$ [44, 42] or they identify a small dominating set [11] of $G$. Another related problem is of dispersion in which $k\leq n$ agents are positioned on $k$ different nodes of $G$, see [34] and the references therein. A solution to the dispersion problem guarantees that $k$ agents are positioned on $k$ different nodes; which is a requirement for the triangle counting problem defined in this paper. Exploration problem on graphs using mobile agents refers to solving a graph analytic task using one or more agents [16].
In this work, we consider triangle counting in a simple, undirected, anonymous graph using mobile agents. The motivation stems from scenarios like private networks in the military or sensor networks in inaccessible terrain where direct access to the network is obstructed, but small battery-powered agents can navigate to learn network structures and their properties for overall network management. Prominent use of agents in network exploration can be seen in areas such as underwater navigation [15], network-centric warfare in military systems [36], modelling social network [56], studying social epidemiology [19] etc.
1.1 Our Contributions
We first enumerate the triangles in the graph $G$ and apply the triangle counting methodology to $3$ applications: (i) Truss Decomposition, (ii) Triangle Centrality, and (iii) Local Clustering Coefficient. Let $G$ be an $n$-node undirected and connected graph with the maximum degree $\Delta$ and diameter $D$. Let $n$ mobile agents with distinct IDs in the range $[0,n^{c}]$ ($c$ being an arbitrary constant) with the highest agent ID $\lambda\in[0,n^{c}]$, be placed at each of the $n$ nodes of $G$ in a dispersed initial configuration. Then, we solve the following problems.
  1. 1.
    Triangle Counting
    • $\bullet$
      Each agent $r_{i}$ can calculate and output the number of triangles adjacent to the vertex where $r_{i}$ is placed on, in $O(\Delta\log\lambda)$ rounds.
    • $\bullet$
      Each agent $r_{i}$ can calculate the number of triangles based on each of its adjacent edges in $O(\Delta\log\lambda)$ rounds.
    • $\bullet$
      Each agent $r_{i}$ can calculate the number of triangles in $G$ in $O(D\Delta\log\lambda)$ rounds.
    The memory requirement for each agent is $O(\Delta\log n)$ bits.
  2. 2.
    Truss Decomposition - The Truss Decomposition Problem for $G$ can be solved by the mobile agents in $O(m\Delta D\log\lambda)$ rounds with $O(\Delta\log n)$ bits of memory per agent.
  3. 3.
    Triangle Centrality - The Triangle Centrality of each node $v\in G$ can be calculated in $O(\Delta\log\lambda)$ rounds if $\mathbf{T(G)}$ is known and in $O(D\Delta\log\lambda)$ rounds, if $\mathbf{T}(G)$ is unknown. $\mathbf{T}(G)$ is the total triangle count of the graph $G$.
  4. 4.
    Local Clustering Coefficient - The Local Clustering Coefficient of each node $v\in G$, i.e., $LCC(v)$ can be calculated by the $n$ agents in $O(\Delta\log\lambda)$ rounds.
The notations are from Table 1.
1.2 Possible Applications
Triangle Counting and Truss Decomposition, in general, have various applications in community detection, social network analysis, biological networks, recommendation systems etc. Mobile agents on the other hand, equipped with truss decomposition capabilities, post dispersion, can dynamically identify and navigate through structurally significant regions of the graph. For example, an unmanned mobile agent can help prioritize areas within a large area that require prompt rescue and can carry rescue materials to difficult-to-access terrain. Similarly, robots carrying different agricultural materials (like fertilizers or pesticides) can dynamically arrange themselves and focus on specific areas that exhibit particular structural characteristics relevant to crop health. Truss decomposition can contribute to the autonomy of vehicles by helping them navigate efficiently through road networks or complex urban environments or helping them identify core areas in a city. In general, mobile robots can prioritize exploring areas with high truss density, indicating regions of potential interest or connectivity within the environment. In addition, its ability to modify itself in challenging scenarios improves fault tolerance and helps maintain the connectivity of the network.
2 Related Work
2.1 Triangle Counting
Triangle counting is a well-studied graph mining problem in both sequential and parallel settings.
2.1.1 Sequential Algorithms:
In 1985, Chiba and Nishizeki [13] proposed an algorithm for counting all triangles in a simple graph by computing the intersection of the neighborhoods of the adjacent vertices with time complexity $O(m^{\frac{3}{2}})$ where $m$ is the number of edges in the graph. Other sequential algorithms for triangle counting are based on vertex-iterator [47] and edge-iterator [26]. In vertex iterator based technique, it iterates over each vertex $v$ of the graph and intersects the adjacency list of each pair of the neighbors of $v$. In edge-iterator based technique, it iterates over each edge and intersects the adjacency list of its two endpoints.
2.1.2 Parallel Algorithms:
In general, a number of distributed and parallel algorithms for triangle enumeration have been proposed for various models (distributed memory, shared memory, multi-core machines, message passing interface (MPI) etc.). In [3], the authors implemented an MPI-based distributed memory parallel algorithm, called PATRIC, for counting triangles in massive networks. Shun et al. in [48], designed multi-core parallel algorithms for exact, as well as approximate, triangle counting and other triangle computations that scale to billions of nodes and edges. In  [4] presented two efficient MPI-based distributed memory parallel algorithms for counting exact number triangles in big graphs. They achieved a faster algorithm using overlapping partitioning and efficient load balancing schemes while a space-efficient one by dividing the network into non-overlapping partitions. Ghosh et al. in [22] presented a simple MPI-based graph triangle counting method for shared and distributed-memory systems, which assumes a vertex-based underlying graph distribution called TriC. It was later improved in [21]. A detailed account of related works on triangle enumeration for various model set-ups may be found in [52, 6, 5, 51, 35, 47, 23, 50, 41].
2.2 Truss Decomposition
In [14], Cohen introduced and visualized truss as a relaxation to cliques in graphs and defined it to be a non-trivial, one-component sub-graph such that each edge is reinforced by at least $k-2$ pairs (for a $k$-truss) of edges making a triangle with that edge. He established the existence of polynomial time algorithms for identifying $k$-trusses within a given graph. Since then, trusses have become the subject of extensive investigation the problem has been studied in various models  [24, 12, 25, 29]. In general, truss decomposition is explored in two main contexts: a serialized version, primarily suitable for small to medium-sized graphs, and a parallel version designed for handling larger graphs. In [53], Yang et al. provided an improved serialized in-memory algorithm for computing $k-$truss in graphs of moderate sizes. In particular, their algorithm computed the $k-$trusses for all $k\geq 3$ in $O(m^{1.5})$ time using $O(m+n)$ memory space. They further proposed two I/O efficient algorithms for handling massive networks which are ill-equipped for the main memory of a single machine. To tackle the large memory requirement for massive graphs, several parallel implementations for truss decomposition were introduced. In [28], the authors implemented a parallel version of the serialized algorithm in [53] and used data structures adaptive to concurrent updates instead of hash tables. Sariyuce et al. in [46] employed the iterative h-index computation, formulated by [38], for nucleus decomposition and proved its convergence bounds. The truss decomposition problem is a particular case of the general nucleus decomposition problem. They developed parallel algorithms for both synchronous and asynchronous versions of the nucleus decomposition problem. In the synchronous version, the computations for specific iterations are made over a single snapshot of current values whereas in the asynchronous version, the latest values of each variable is used instantly in the current computation. Voegele et al. in [52] proposed a parallel graph-centric $k-$ truss decomposition and established the relation between a $k-$ truss and a $k-$core. When a $k-$ core computation proceeds $k-$ truss, a lot of edges from the graph may be waived off from being processed since a $k-$ truss is always contained in a $k-1$ core. Jian Wuet al. in [55], engineered both the serialized and parallel algorithms of  [53] and  [46] to reduce their memory usage by optimizing the underlying data structures and by using WebGraph. In [20], the authors implemented the truss decomposition on probabilistic graphs. They proposed an algorithm based on h-index updating and obtained an upper bound on the number of iterations for convergence. The efficiency of most of the algorithms is described with the aid of extensive experimental results.
Agent-based Computations on Graphs: In [49], Sudo et al. considered the exploration problem with a single agent in undirected graphs. Starting from an arbitrary node, the agent has to explore all the nodes and edges in the graph and return to the starting node. The authors used a whiteboard model, that reduced the memory requirement per robot. In  [30], the authors explored the graph exploration problem based on Depth-First-Search and studied the trade-off between node-memory vs robot-memory. In [17], Dereniowski et al. proposed an algorithm for collective graph exploration with a team of $k$ ($k$ being polynomial size) agents in $O(D)$ time. They also obtained almost tight bounds on the asymptotic relation between exploration time and team size, for large $k$, in both the local and the global communication model. Further results on agent-based graph exploration and dispersion (spreading agents across the graph so that there are at-most one agent at each node) can be found in several recent papers [31, 32, 33, 39, 40, 43, 10].
3 Model and Problem Definitions
Graph: The underlying graph $G(V,E)$ is connected, undirected, unweighted and anonymous with $|V|=n$ nodes and $|E|=m$ edges. The nodes of $G$ do not have any distinguishing identifiers or labels. The nodes do not possess any memory and hence cannot store any information. The degree of a node $v\in V$ is denoted by $\delta(v)$ and the maximum degree of $G$ is $\Delta$. Edges incident on $v$ are locally labelled using port numbers in the range $[0,\delta(v)-1]$. A single edge connecting two nodes receives independent port numbering at the two ends. The edges of the graph serve as routes through which the agents can commute. Any number of agents can travel through an edge at any given time.
Mobile Agents (or Mobile Robots): We have a collection of $n$ agents =$\mathcal{R}=\{r_{1},r_{2},...,r_{n}\}$ residing on the nodes of the graph in such a way that each node is occupied by a distinct ID agent at the start (known as dispersed configuration in literature). Each agent has a unique ID in the range $[0,n^{c}]$ ($c$ is an arbitrary constant) and has $O(\Delta\cdot\log(n))$ bits to store information. An agent therefore cannot store the whole graph structure information within its limited memory. An agent retains its memory as long as needed and it can be updated as required. Two or more agents can be present (co-located) at a node or pass through an edge in $G$. However, an agent is not allowed to stay on an edge. An agent can recognise the port number through which it has entered and exited a node. The agents do not have any visibility beyond their (current) location at a node. An agent at node $v$ can only see the adjacent ports (connecting to edges) at $v$. Only the collocated agents at a node can sense each other and exchange information. An agent can transfer all the information stored in its memory in a single round. We interchangeably use the terms robots and agents.
Communication Model: We consider a synchronous system where the agents are synchronised to a common clock. As mentioned earlier, we consider the local communication model where only co-located agents (i.e., agents at the same node) can communicate among themselves.
Time Cycle: Each agent $r_{i}$, on activation, performs a $Communicate-Compute-Move$ $(CCM)$ cycle as follows.
    Communicate: $r_{i}$ may communicate with other agents present at the same node as itself.
    Compute: Based on the gathered information and subsequent computations, $r_{i}$ may perform all manner of computations within the bounds of its memory.
    Move: $r_{i}$ may move to a neighbouring node using the computed exit port.
An agent can perform the $CCM$ task in one time unit, called round. The time complexity of an algorithm is the number of rounds required to achieve the goal. The memory complexity is the number of bits required by each agent to execute the algorithm.
The following Table 1 provides a quick reference of the notations used throughout the paper.
Symbols Meaning
$G(V,E)$ Graph with edge-set $E$ and node-set $V$
$n,m$ Number of nodes and edges of G respectively
$\mathcal{R}$ Collection of $n$ mobile robots $\{r_{1},r_{2},\dots,r_{n}\}$
$\lambda$ ID of the robot with maximum ID
$\delta(v)$ Degree of the node $v$
$\Delta$ Highest degree of node in $G$
$D$ Diameter of $G$
$T_{k}$ $k$-truss of $G$
$T(v)$ Number of triangles with $v$ as vertex (node)
$T(G)$ Total triangle count of $G$
$N(v)$$\{u:(u,v)\in E(G)\}$
$N_{\mathbf{T}}(v)$$\{u\in N(v):N(u)\cap N(v)\neq\phi\}$
$N^{+}_{\mathbf{T}}(v)$$\{v\}\cup N_{\mathbf{T}}(v)$
$LCC(v)$ Local Clustering Coefficient of node $v$
Table 1: Notations used in the paper.
3.1 Problem Statements
Problem 1.
(Triangle Counting using Mobile Agents) Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ agents, each of which initially placed distinctly at each node of $G$. We solve the following problems.
  1. 1.
    Node-Based Triangle Counting: To count the number of triangles with a given node as a vertex.
  2. 2.
    Edge-Based Triangle Counting: To count the number of triangles based on a given edge.
  3. 3.
    Total Triangle Counting: To count the total number of triangles in the graph $G$.
Problem 2.
(Truss Decomposition using Mobile Agents) Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of $n$ agents, each of which are initially placed distinctly at each node of $G$. The $n$ autonomous agents coordinate among themselves to solve the Truss Decomposition Problem.
Problem 3.
(Triangle Centrality using Mobile Agents) Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of $n$ agents, each of which initially placed distinctly at each node of $G$. The $n$ autonomous agents coordinate among themselves to compute the Triangle Centrality $TC(v)$ for each node $v\in G$.
Problem 4.
(Local Clustering Coefficient using Mobile Agents) Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of $n$ agents, each of which initially placed distinctly at each node of $G$. The $n$ autonomous agents coordinate among themselves to compute the Local Clustering Coefficient, $LLC(v)$ for each node $v\in G$.
In this paper, we study the above problems from a theoretical perspective and aim to solve them while minimizing both time and memory-per-agent as much as possible.
4 Triangle Counting via Mobile Agents
In this section, we develop algorithms for $n$ mobile agents that are initially dispersed among the $n$ nodes of the graph $G$ to enumerate the number of triangles in $G$. Since the nodes themselves are memory-less and indistinguishable, the algorithm relies on the memory and IDs of the mobile agents that reside on the nodes of the graph. Also, since the agents cannot communicate among themselves (unless they are at the same node), synchronising the movement of the agents is another challenge.
In our algorithm, the agents (technically representing the nodes they are sitting at) first scan their neighbourhood. Once all the information about the neighbourhood is collected, the agents now count the number of common neighbourhoods between two adjacent agents. After each agent receives that count, the sum of each such count stored at each agent is evaluated, which when divided by three gives us the number of triangles in $G$. The algorithm runs in three phases. In the first phase, the agents learn their neighbours. In the second phase, the agents check the number of common neighbours with each of its adjacent agents. In this phase, each agent $r_{i}$ also counts the number of local triangles with $r_{i}$ as a vertex and the number of triangles with $(r_{i},r_{j})$ as an edge, where $r_{j}$ is an adjacent agent to $r_{i}$. In the third and last phase, each agent collects the local triangle count from every other agent and counts the number of triangles in $G$. We now explain each phase in detail, below.
4.1 Phase 1: Know the Neighbourhood
Each node is occupied by a distinct (via their IDs) agent before the start of the algorithm and we represent each node of $G$ by its stationed agent. The algorithm starts with Phase 1, where each agent discovers and records its neighbour. As the IDs of the agents are unknown, agents cannot synchronously start scanning the neighbourhood. With each agent executing the same algorithm, agents may not find other agents at their exact place when they move. Therefore, we need to ensure that each gets to record the correct neighbour set of it. To do this we exploit the ID bits of the agents. We use the fact the IDs of the agents are distinct. To this end, we state the following lemma.
Lemma 1.
Let $r_{i}$ and $r_{j}$ be two distinct agents in $\mathcal{R}$, with their IDs field being $r_{i}.ID$ and $r_{j}.ID$ respectively. Then, there exists at least one dissimilar bit in $r_{i}.ID$ and $r_{j}.ID$ with one being 00 and the other being $1$.
Let $\lambda$ denote the largest ID among all the $n$ agents. Therefore, the agents use a $\log\lambda$ bit field to store the IDs. Now, to list the neighbouring agents, an agent $r_{i}$ stationed at a node $u$ does the following. Here, $\delta(u)$ denotes the degree of a node $u$ and $\Delta$ denotes the highest degree of a node in $G$.
  1. 1.
    For $\log\lambda$ rounds $r_{i}$ executes the following.
    1. (a)
      $r_{i}$ checks the current ID bit in its ID from the right. (At the start, the current ID bit is the rightmost bit in the ID field).
    2. (b)
      In the next $2\Delta$ rounds, $r_{i}$ chooses to do one of the following two:
      • $\bullet$
        If the current ID bit is 00, $r_{i}$ waits at its own node for $2\Delta$ rounds.
      • $\bullet$
        If the current ID bit is $1$, $r_{i}$ visits each of its neighbour and back using port number 00 till port $\delta(u)-1$. When $r_{i}$ meets a new agent $r_{k}$, it checks $r_{k}$’s current ID bit. If the current ID bit of $r_{k}$ is 00, it adds $r_{k}$ and records it to $r_{i}$’s neighbours list. (This is to ensure that $r_{k}$ is the original neighbour of $r_{i}$ and $r_{k}$ is not an exploratory agent from a different neighbouring node). The agent $r_{k}$ also simultaneously records $r_{i}$ as its neighbour as well. The agents ignore any agent that has already been registered.
    3. (c)
      After the $2\Delta$ rounds have elapsed, the next left bit in the ID field becomes the current ID bit. If no more bits are remaining and $\log\lambda$ rounds have not been completed (implying that $r_{i}$ has a smaller ID length than $\log\lambda$ bits), $r_{i}$ assumes the current bit as 00 and stays back at its own node for the rest part of the algorithm.
After the completion of Phase 1 of our algorithm, it is guaranteed that each agent correctly records all its neighbouring agents in $O(\Delta\cdot\log\lambda)$ rounds, which we prove in the following lemma.
Lemma 2.
Each agent $r_{i}$ correctly records the exhaustive list of its neighbour agents in $O(\Delta\cdot\log\lambda)$ rounds after the completion of Phase 1.
Proof.
Let $r_{j}$ (placed at node $v$) be a neighbour of $r_{i}$ (at node $u$). Then there exists a port joining $u$ to $v$ having port number between $[0,\delta(u)-1]$. Since $r_{i}$ and $r_{j}$ have distinct IDs, there exists a bit (say the $p^{th}$ bit from the right) in $r_{i}.ID$ and $r_{j}.ID$, which are different from one another (one bit being 00, the other being $1$)[Lemma 1]. Without the loss of generality, let us assume that the $p^{th}$ bit of $r_{i}$ is $1$ and the $p^{th}$ bit of $r_{i}$ is 00. Therefore, when the $p^{th}$ bit becomes the current ID bit, $r_{i}$ starts exploring its neighbours one by one during which it finds the agent $r_{j}$ now stationary at $v$ with current ID bit 00. Therefore $r_{i}$ records $r_{j}$ as its neighbour (simultaneously $r_{j}$ also registers $r_{i}$ in its neighbour list). In a similar way, $r_{i}$ records all of its neighbours exhaustively either by visiting a stationary agent in its neighbourhood or by meeting an exploratory agent from its neighbour. Phase 1 runs for $\log\lambda$ rounds with each round consisting of $O(\Delta)$ sub-rounds to allow the agents to visit each neighbour. Therefore, Phase 1 completes in $O(\Delta\cdot\log\lambda)$ rounds. ∎
Therefore, with the end of Phase 1, each agent has now enlisted its neighbouring agents in its memory. The algorithm now moves to Phase 2 where each agent $r_{i}$ counts the number of triangles with $r_{i}$ as one of its vertex (node).
4.2 Phase 2: Local Triangle Counting
In phase two, each agent, now equipped with the list of its neighbouring agents, visits each of its neighbours once again, in exactly the similar fashion described in Phase 1. Whenever the agent $r_{i}$ meets its neighbour $r_{j}$, they communicate to find out the number of common neighbours they both have. Here as the nodes are anonymous, they are identified using the mobile agents that reside on the nodes. This communication is also used the update the following variables of $r_{i}$:
  1. 1.
    $r_{i}.edge(r_{j}):$ A variable which stores the count of the number of common neighbours of $r_{i}$ and $r_{j}$ which represents the number of triangles based on the edge containing $(r_{i},r_{j})$. $r_{i}$ has $\Delta$ such variables $r_{i}.edge(x)$, where $x$ represents another agent that has an edge with $r_{i}$. Each of these variables are initially set to 00. A variable $r_{i}.edge(x)$ remains 00, if there are no triangles with $(r_{i},x)$ as an edge.
  2. 2.
    $r_{i}.local\_sum:$ Initially set to 00, adds up the counts of the number of common neighbours for each distinct neighbour of $r_{i}$. As $r_{i}$ finds its neighbours one by one, it cumulatively adds up the count of the number of common neighbours with each of its neighbours. Mathematically, $r_{i}.local\_sum=\Sigma_{j}r_{i}.edge(r_{j})$, where the sum runs over every neighbouring agent $r_{j}$ of $r_{i}$.
In the given window of $\log\lambda$ rounds (each round containing a $\Delta$ sub-round), as the agents communicate with their neighbours, the variables $r_{i}.edge(\cdot)$ are updated. At the end of this phase, $r_{i}$ also builds on the count of the variable $r_{i}.local\_sum=\Sigma_{j}r_{i}.edge(r_{j})$. To get the correct number of triangles with $r_{i}$ as one of the vertex (local triangle counting), we divide $r_{i}.local\_sum$ by $2$. We state the reason for the same in the following lemma.
Lemma 3.
The number of triangles with agent $r_{i}$ as a vertex is given by $\frac{1}{2}(r_{i}.local\_sum)$.
Proof.
First, we see that if the agents $r_{i}$ and $r_{j}$ have a common neighbour $r_{k}$, then, $(r_{i},r_{j},r_{k})$ form a triangle with $r_{i}$ as one of the vertex. Now, the triangle $(r_{i},r_{j},r_{k})$ has been tallied once while counting the common neighbour of $r_{i}$ with $r_{j}$ and re-counted again as a common neighbour of $r_{i}$ with $r_{k}$. Therefore, each $(r_{i},r_{j},r_{k})$ with $r_{i}$ as vertex, is counted twice, once through $r_{j}$ and other time through $r_{k}$. So, the number of triangles with agent $r_{i}$ as a vertex is exactly half of $r_{i}.local\_sum$. ∎
Therefore at the end of Phase 2, each agent $r_{i}$ gets a local count of the number of triangles with $r_{i}$ as a vertex and a list of the number of triangles that are based on the edges adjacent to $r_{i}$. In the next phase, the algorithm finds the number of triangles of $G$ using the counts generated in this phase.
4.3 Phase 3: Counting the Number of Triangles in $G$
To find the number of triangles in $G$, we need to take into account the local triangle counts of each agent. We first need to accumulate the $r_{i}.local\_sum$ values of each of the $n$ agents and calculate the triangles of $G$ from there. However, due to the absence of a $leader$ agent and the difficulty of synchronising the movement of the agents, gathering the $r_{i}.local\_sum$ is not straightforward. The high-level idea is to communicate repeatedly with the neighbours and continually gather up the values of $local\_sum$ of each agent hop by hop till each agent gets every $local\_sum$ value.
As described in the previous case, the agents can assuredly communicate with all their neighbours in $\Delta\cdot\log\lambda$ rounds. Inside the first $\Delta\cdot\log\lambda$ round of this phase, the agents communicate with their neighbours and exchange the variable $local\_sum$. An agent $r_{i}$, along with its own $r_{i}.local\_sum$, collects $r_{j}.local\_sum$ values for each of its neighbour $r_{j}$ and stores it in the memory. Within the next $\Delta\cdot\log\lambda$ rounds, $r_{i}$ meets its neighbour agents $r_{j}$ again to check if they have collected any new $local\_sum$ values (possibly from their own neighbour). $r_{i}$ again stores any new $local\_sum$ record (of its neighbour’s neighbour) that it receives through its neighbour $r_{j}$ in the current phase. By the end of the second phase, $r_{i}$ has the information of the $local\_sum$ of all the agents that are at a distance of $2-hops$ from it. Now, with the completion of $D$ (the diameter of the graph $G$) such $\Delta\cdot\log\lambda$ rounds, each agent $r_{i}$ now has the $local\_sum$ record of every agent in $G$. The number of triangles in $G$ is finally calculated by each agent by summing up the $n$ $local\_sum$ values that they have gathered and dividing it by $6$.
Lemma 4.
Let $r_{i}$ and $r_{k}$ be two agents at a distance of $d$ hops from each other. Then, the value of $r_{k}.local\_sum$ can be communicated to $r_{j}$ within $d\cdot O(\Delta\cdot\log\lambda)$ rounds.
Proof.
Since $r_{i}$ and $r_{k}$ are at a distance of $d$ hops from each other, there exists a sequence of agents (nodes) $(r_{i},r_{i_{1}},r_{i_{2}},\dots,r_{i_{d-1}},r_{k})$ from $r_{i}$ to $r_{k}$. In the first $O(\Delta\cdot\log\lambda)$ rounds, $r_{i}$ can communicate with $r_{i_{1}}$ to get the value of $r_{i_{1}}.local\_sum$. In the meanwhile, $r_{i_{1}}$ also collects the value of $r_{i_{2}}.local\_sum$ in the same $O(\Delta\cdot\log\lambda)$ round. Therefore, in the second sub-phase of $O(\Delta\cdot\log\lambda)$ rounds, when $r_{i}$ communicates with $r_{i_{1}}$ again, $r_{i}$ receives the value of $r_{i_{2}}.local\_sum$ through $r_{i_{1}}$. Continuing in a similar way, $r_{i}$ is guaranteed to receive the value of $r_{k}.local\_sum$ from $r_{k}$ at a distance of $d$ hops through the agents $r_{i_{1}},r_{i_{2}},\dots,r_{i_{d-1}}$ by the end of $d\cdot O(\Delta\cdot\log\lambda)$ rounds. ∎
Let $D$ denote the diameter of the graph $G$. Since any two agents in $G$ are located at a distance of at most $D$ hops, the following lemma follows from Lemma 4.
Lemma 5.
After the end of Phase 3, which takes $O(D\cdot\Delta\cdot\log\lambda)$ rounds, each agent $r_{i}$ has the complete record of $local\_sum$ values of every agent in $G$. Here, $D$ is the diameter of the graph $G$.
If the diameter $D$ of the graph is unknown, $D$ could be replaced by $n$.
Lemma 6.
The number of triangles in graph $G$ is given by $\frac{1}{6}\cdot\Sigma_{i}(r_{i}.local\_sum)$, where the sum runs over all the $n$ agents of $G$.
Proof.
The number of triangles with agent $r_{i}$ as a vertex is given by $\frac{1}{2}\cdot r_{i}.local\_sum$. Now since each triangle is counted once for every vertex it has, the total number of triangles in $G$ is given by $\frac{1}{3}\cdot\Sigma_{i}\frac{1}{2}\cdot(r_{i}.local\_sum)$. ∎
Notation: We shall denote the number of triangle containing the vertex(node) $v$ and the total number of triangles in $G$ with $\mathbf{T}(v)$ and $\mathbf{T}(G)$, respectively. For a node $v$ with a robot $r_{i}$ on it, we use $r_{i}$ and $v$ interchangeably to denote the node.
At the end of this phase, each agent has the value of the number of triangles in $G$ i.e., $\mathbf{T}(G)$. In the following theorem, we assemble the list of results we got during the 3 phases.
Theorem 4.1.
Let $G$ be an $n$ node arbitrary, simple, connected graph with a maximum degree $\Delta$ and diameter $D$. Let $n$ mobile agents with distinct IDs in the range $[0,n^{c}]$ with the highest agent ID $\lambda\in[0,n^{c}]$, where $c$ is constant, be placed at each of the $n$ nodes of $G$ in a dispersed initial configuration. Then,
  1. 1.
    Each agent $r_{i}$ can calculate the number of triangles with $r_{i}$ as a vertex i.e., $\mathbf{T}(r_{i})$ in $O(\Delta\cdot\log\lambda)$ rounds.
  2. 2.
    Each agent $r_{i}$ can calculate the number of triangles based on each of its adjacent edges in $O(\Delta\cdot\log\lambda)$ rounds.
  3. 3.
    Each agent $r_{i}$ can calculate the number of triangles in $G$, $\mathbf{T}(G)$, in $O(D\cdot\Delta\cdot\log\lambda)$ rounds.
5 Applications: Truss Decomposition, Triangle Centrality and Local Clustering Coefficient
In this section, we show some novel applications of our triangle-counting algorithm via mobile agents in an anonymous graph.
5.1 Truss Decomposition
We first use the techniques from Section 4 to identify a $k$-truss sub-graph of a given graph $G$ using mobile agents, when it exists. A sub-graph $T_{k}$ of $G$ is called a $k$-truss if every edge of $T_{k}$ is a part of at least $k-2$ triangles i.e, each edge in $T_{k}\subset G$ is supported by at least $k-2$ triangles in $T_{k}$.
In our algorithm, we use the popular truss decomposition method to first calculate, what is called, $trussness$ value for each edge. The $trussness$ values are then used to construct the $k$-truss sub-graph for any value of $k$.
Before we explain our algorithms in detail, we formally define $support$, $k$-truss and $trussness$.
Definition 1 ($support$).
For a given graph $G(V,E)$, the $support$ of an edge $e\in E$ is the number of $triangles$ in $G$ that contain $e$.
Definition 2 ($k$-truss).
[…]
Definition 3 ($trussness$).
The $trussness$ of an edge $e$, is defined as the maximum $k$ such that $e$ belongs to $T_{k}$ but does not belong to $T_{k+1}$.
Algorithm for Truss Decomposition: Let us now propose an algorithm for mobile agents to evaluate $trussness$ for each edge of the graph $G(V,E)$. The algorithm is based on the parallel truss decomposition described in [55]. The $trussness$ values determine a partition (thus an equivalence relation) on $E$, where each class have the edges of $G$ with the same $trussness$ value. Let $\mathbf{t_{k}}$ denote the equivalence class of edges having $trussness=t_{k}$. To find the $k$-truss, $T_{k}$ we construct a sub-graph of $G$ with edges from the equivalence classes $\mathbf{t_{k}},\mathbf{t_{k+1}},\dots,\mathbf{t_{max}}$, where $t_{max}$ is the maximum $trussness$ of any edge in $G$. Mathematically, $T_{k}=\cup_{i=k}^{t_{max}}\mathbf{t_{i}}$. Therefore by computing the trussness for each edge in $G(V,E)$, we obtain a partition (equivalence classes) of $E$, thereby obtaining the $k$-trusses of $G$ for any $k$ by taking the union of the equivalent classes. Therefore, the $k$-truss decomposition of a graph is equivalent to computing the $trussness$ of each edge in the graph.
$trussness$ values of each edge in $G$ can be computed efficiently using existing serial and parallel algorithms. The serial truss decomposition algorithm first calculates $support$ for each edge and iteratively removes a single edge each time until all the edges of the graph are removed. Each time, the edges are sorted in ascending order and the edge with the lowest $support$ is removed. The removed edge (say $e$) keeps its final $support+2$ as its $trussness$ value. Once the $e$ is removed, the $support$ value of the edges that formed the triangle with $e$ is re-evaluated. The remaining edges are once again sorted in order before running another pass. The need to sort the edges in order of $support$ makes the algorithm inherently sequential.
In the parallel version of the algorithm, the sorting condition is relaxed with the use of $h-index$ updating. For a set of integers $S$, the $h-index$ of $S$ is defined as the largest number $h$ such that there are at least $h$ elements in $S$ that are equal to and greater than $h$. The $trussness$ of an edge is related to $h-index$, for, the $trussness$ of an edge $e$ can be thought of as the largest $k$ such that it is contained in at least $k$ triangles whose edges have a $trussness$ value of at least $k$. In the algorithm, each edge $e$ in $G$ is initialized to its $support$ as the first approximation to its $trussness$. Now, the $support$ values of all triangles with $e$ as an edge are stored in a set $L$ and its $h-index$ is computed. At each iteration, the $h-index$ of $e$ is updated to the smallest of its current value and the $h-index$ of $L$. The algorithm iteratively updates an edge’s $h-index$ by computing the $h-index$ of all edges that support it, until achieving convergence when no updates would happen. The final $h-index$ of each edge before, no further updates happen, provide the $trusness$ value for each edge.
However, mobile agents bring unique challenges. The agents do not have any global knowledge of the topology of the network and can only communicate with other agents if they are located in the same node (or they are in sufficient proximity). Further, the movement of the agents, although synchronous, does not have any centralized control. These limitations pose unique challenges in constructing algorithms for truss decomposition. We, therefore, engineer the existing parallel version [55] of the truss decomposition algorithm to adapt to our mobile agent model.
To start with, each of the $n$ agents, $r_{i}$ has the following variables:
  1. 1.
    $r_{i}.ID$ stores the ID string of the agent $r_{i}$. As per our assumption, agents have ID in the range $[0,n^{c}]$, where $c$ is an arbitrary but fixed constant, so $r_{i}.ID$ requires $O(\log n)$ bits of storage space.
  2. 2.
    $r_{i}.edge\_set$ to store the edges incident on a particular node containing $r_{i}$. $r_{i}$ stores only edges $(r_{i},r_{j})$, where $r_{j}>r_{i}$. $edge\_set$ requires $O(\Delta)$ bit of memory, where $\Delta$ is the highest degree of a node in $G$. The ordering of edges ensures that each edge becomes associated with exactly one agent.
  3. 3.
    $h$ a variable associated with each edge in $edge\_set$; initially stores the $support$ of each edge. $h$ is next updated according to the $h-index$ values as the algorithm progresses, as described earlier. The final $h$ values provide the $trussness$ for a particular edge.
  4. 4.
    $L$ also associated with each edge in $edge\_set$. $L$ is used to store the values $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$ for for each triangle $(e,e^{\prime},e^{\prime\prime})$ containing $e$.
  5. 5.
    $N$ is used to temporarily store the edges $e^{\prime},e^{\prime\prime}$ for each triangle $(e,e^{\prime},e^{\prime\prime})$ containing $e$.
  6. 6.
    $scheduled$ to identify the edges in the $edge\_set$ that need further updating in consequent rounds of the algorithm. Initially, every edge in $edge\_set$ is scheduled ($true$).
  7. 7.
    $r_{i}.change$ initially assigned 00. If no more edges in $r_{i}.edge\_set$ are scheduled, $r_{i}.change$ becomes $1$.
Our algorithm for mobile agents runs in Phases as described below, in detail.
5.1.1 Phase 1 (Registering neighbour agents, adjacent edges and port numbers)
With each node of $G$ hosting a distinct agent, each agent first visits and registers its list of neighbours. Due to a lack of movement synchronisation, the agents follow the exact protocol as described in Phase 1 of Section 4 to register the neighbours, the ports which lead to the respective neighbours and the respective edges. To avoid inconsistency, the agents maintain the record of the edges in a specific order. An agent $r_{i}$ only stores the record of an edge $(r_{i},r_{j})$ only if $r_{j}$ is a neighbour of $r_{i}$ and $r_{j}.ID>r_{i}.ID$. This gives a total ordering on the edges. This phase takes $O(\Delta\cdot\log\lambda)$ rounds to complete. Now the agents begin the next phase of calculating $support$ for each edge. With the execution of Phase 1, $edge\_set$ gets recorded.
5.1.2 Phase 2 ($support$ calculation and initial $h$ initialization)
The agents now calculate the support for each edge. To do this, the agents use a similar methodology to the one used in Phase 2 of Section 4. The agents go to each of their neighbours once again. Let us assume, at some point of time during the algorithm, an agent $r_{i}$ meets its neighbour $r_{j}$. $r_{i}$ then checks the number of common neighbours with $r_{j}$ and records the count into $r_{i}.support(r_{i},r_{j})$ (we have assumed that $r_{j}.ID>r_{i}.ID$; otherwise if $r_{j}.ID<r_{i}.ID$, then the edge $(r_{i},r_{j})$ (simultaneously, $r_{i}.support(r_{i},r_{j})$) does not register in $r_{i}$’s records and in such case, the agent $r_{i}$ does nothing. However, in such case, the edge does register in $r_{j}$’s record, and $r_{j}$ then does the needful in that respect.). Support calculation takes $O(\Delta\log\lambda)$ rounds, after which, each agent receives its $support$ number. With the execution of this, $h$ value for each edge in $r_{i}.edge\_set$ gets initialized to the support of the particular edge.
5.1.3 Phase 3 (Iterative $h-index$ up-gradation)
With each edge receiving its initial $support$ value, the algorithm now updates $h$ for each edge through multiple iterations of this phase. The phase iterates till the $change$ value in any of the agents $r_{i}$ remains 00 and terminates once the value of $change$ in each and every agent becomes $1$. In each running phase, each agent $r_{i}$ does the following in parallel.
  1. 1.
    Checks if any edge is scheduled for further up-gradation in its $edge\_set$. If no, set $r_{i}.change\leftarrow 1$ and stops executing the algorithm but remains active. $r_{i}$ provides any information when required to other agents.
  2. 2.
    Resets the sets $L$ and $N$.
  3. 3.
    For each edge $e$ in $edge\_set$ which is scheduled for up-gradation, $L$ stores $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$ for each triangle $(e,e^{\prime},e^{\prime\prime})$ in $G$. To obtain $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$, $r_{i}$ must communicate with its neighbouring agent. For example, consider the edge $(r_{i},r_{j})$ with $r_{i}<r_{j}$, stored in $r_{i}.edge\_set$. Also, consider another agent $r_{p}(r_{p}<r_{i}<r_{j})$ stationed on another node such that $(r_{i},r_{j},r_{p})$ forms a triangle. Let $e=(r_{i},r_{j}),e^{\prime}=(r_{p},r_{i}),e^{\prime\prime}=(r_{p},r_{j})$. Now for $r_{i}$ to calculate $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$, $r_{i}$ must visit $r_{p}$. Since the movement of the agents is not centralised, $r_{i}$ must execute "Know the Neighbourhood" (Phase 1, Sec. 4.1) which takes additional $O(\Delta\log\lambda)$ rounds, in order to guarantee that $r_{i}$ meets $r_{p}$. Along with the set $L$, the set $N$ is also updated. Let $e=(r_{i},r_{j})$. For each triangle $(e,e^{\prime},e^{\prime\prime})$ the agents add $e^{\prime},e^{\prime\prime}$ to their respective $N$ sets. Now, since each edge, $e^{\prime}$ or $e^{\prime\prime}$ is associated with a unique agent, an invocation of "Know the Neighbourhood" (Phase 1, Sec. 4.1) is required to inform the correct agent associated with a particular edge so that the agent can schedule the particular edge in the next step if required.
  4. 4.
    $r_{i}$ now computes the $h-index$ for the set $L$. If the calculated $h-index$ is less than $h(e)$, then, $e$ along with the edges $e^{\prime}\in N$ with $h-index(L)<h(e^{\prime})\leq h(e)$ is further scheduled for up-gradation in the next iteration and $h(e)$ is updated to the new $h-index$. Otherwise, $e$ is unscheduled and its current $h$ value gives the $trussness$ at the end of the algorithm (if not updated further during the algorithm).
  5. 5.
    Based on whether there are scheduled edges in $r_{i}.edge\_set$, $r_{i}.change$ updates its value if necessary.
  6. 6.
    Checking for Termination: Each agent, after having evaluated the $change$ value for the current phase, communicates with each of the other agents to check, if the algorithm is ready for termination. The agent sends the $change$ values to every agent in multiple rounds. To communicate, each agent starts by visiting its immediate neighbours. In the first round, each agent sends the $change$ to its neighbour. In the second round, the same change value is propagated to the 2-hop neighbour and so on. By the end of $D$ rounds, each agent receives every changed value from every other agent. Now, the agents calculate the binary product of all the $change$ values received. If the binary product of the $change$ value results in $1$ (all the agents must get the same binary product value), the agents terminate the algorithm. Otherwise, the agents rerun Phase 3 again, till the product becomes $1$.
Lemma 7.
Algorithm 5.1 takes $O(m\Delta D\log\lambda)$ rounds to terminate
Proof.
Phase 1 and Phase 2 takes $O(\Delta\log\lambda)$ rounds, as shown in Lemma 2 and Lemma 3. In each execution of Phase 3, the agent needs to communicate with its neighbour to update its $trussness(h)$ values taking $O(\Delta\log\lambda)$ rounds. Finally, to check for termination, each agent propagates the $change$ value to every other agent, needing $O(D\Delta\log\lambda)$ rounds.
At worst, the algorithm may update the $h-index$ value of one edge during each execution of Phase 3. Also, the up-gradation of the $h-index$ value of an edge is equivalent to the fact that the triangle count ($support)$ of that particular edge (and possibly some other edges) in $G$ has been decreased. Therefore, each time an $h-index$ of an edge is updated (it can only decrease), the graph $G$ loses at least one edge. Therefore, the algorithm may need to execute Phase 3, $m$ times before termination. Therefore Algorithm 5.1 takes $O(\Delta\log\lambda)+m\cdot(O(\Delta\log\lambda)+O(D\Delta\log\lambda))=O(m\Delta D\log\lambda)$ rounds to execute. ∎
Theorem 5.1.
Let $G$ be an $n$-node arbitrary, simple, connected graph with a maximum degree $\Delta$ and diameter $D$. Let $n$ mobile agents with distinct IDs in the range $[0,n^{c}]$ with the highest agent ID $\lambda\in[0,n^{c}]$, where $c$ is constant, be placed at each of the $n$ nodes of $G$ in a dispersed initial configuration. Then, the Truss Decomposition Problem for $G$ can be solved by the mobile agents in $O(m\Delta D\log\lambda)$ rounds with $O(\Delta\cdot\log n)$ bits of memory per agent.
5.2 Triangle Centrality
The concept of Triangle Centrality was introduced in  [9] by Paul Burkhardt. The concept may be useful for finding important vertices in a graph based on the concentration of triangles surrounding each vertex. An important node in triangle centrality is at the centre of many triangles, and therefore it may be in many triangles or none at all. In this section, we employ $n$ mobile agents, each with a distinct ID starting at each distinct node of an arbitrarily connected anonymous graph to compute the triangle centrality for each node of the graph.
Mathematically, Triangle Centrality, $TC(v)$ of a node $v\in G$ is formulated in  [9] as:
$TC(v)=\frac{\frac{1}{3}\sum_{u\in N^{+}_{\mathbf{T}}(v)}\mathbf{T}(u)+\sum_{w\in N(v)\setminus N_{\mathbf{T}}(v)}\mathbf{T}(w)}{T(G)}$
where, where $N(v)$ is the neighborhood set of $v$, $N_{\mathbf{T}}(v)$ is the set of neighbors that are in triangles with $v$, and $N^{+}_{\mathbf{T}}(v)$ is the closed set that includes $v$. $\mathbf{T}(v)$ and $\mathbf{T}(G)$ denote the respective triangle counts based on $v$ and total triangle count in $G$. Mathematically, $N(v)=\{u:(u,v)\in E(G)\}$, $N_{\mathbf{T}}(v)=\{u\in N(v):N(u)\cap N(v)\neq\phi\}$ and $N^{+}_{\mathbf{T}}(v)=\{v\}\cup N_{\mathbf{T}}(v)$. Now, we outline our algorithm that computes $TC(v)$ for a node $v\in G$ using the mobile agents that runs in these $3$ phases:
  1. 1.
    Computing $\mathbf{T}(v)$ and $\mathbf{T}(G)$ We use the triangle counting algorithm for mobile agents described in Section 4 to evaluate the number of triangles involving the node $v$, $\mathbf{T}(v)$ and the total number of triangles in $G$, $\mathbf{T}(G)$. The execution of this phase takes $O(\Delta\log\lambda)+O(D\Delta\log\lambda)=O(D\Delta\log\lambda)$ rounds [Lemma 3,6]. With the completion of this phase, the agents begin the next phase. Note that if the count of the total number of triangles is available apriori, only $T(v)$ needs to evaluated.
  2. 2.
    Computing $N(v)$, $N^{+}_{\mathbf{T}}(v)$ and $N_{\mathbf{T}}(v)$ $N(v)$, the neighbours of $v$ can be recorded in $O(\Delta\log\lambda)$ rounds by the agents (Lemma 2). After we allow the agents to record their neighbourhood in the first $O(\Delta\log\lambda)$ rounds, in the subsequent $O(\Delta\log\lambda)$ rounds, the agents can now meet their neighbours once again, with this time, learning about $N(u)$ for each neighbour agent $u$ of $v$. Now the agent from $v$ communicates with each $u\in N(v)$ to evaluate the set $N_{\mathbf{T}}(v)$. Similarly, the agents can also build the set $N^{+}_{\mathbf{T}}(v)$. Now, as the agent meets with each of its neighbours $u\in N(v)$, it cumulatively evaluates the sums $s=\sum_{u\in N^{+}_{\mathbf{T}}(v)}$ and $x=\sum_{u\in N(v)_{\mathbf{T}}(v)}$ simultaneously. Finally the sum $\sum_{w\in N(v)\setminus N_{\mathbf{T}}(v)}=s-x+\mathbf{T}(v)$ is evaluated. This step takes $O(\Delta\log\lambda)$ rounds.
  3. 3.
    Computing $TC(v)$ With the values of $\sum_{u\in N^{+}_{\mathbf{T}}(v)}\mathbf{T}(u)$,
    $\sum_{w\in N(v)\setminus N_{\mathbf{T}}(v)}\mathbf{T}(w)$ and $T(G)$ have now been obtained, the agent at $v$ can now evaluate
    $TC(v)=\frac{\frac{1}{3}\sum_{u\in N^{+}_{\mathbf{T}}(v)}\mathbf{T}(u)+\sum_{w\in N(v)\setminus N_{\mathbf{T}}(v)}\mathbf{T}(w)}{T(G)}$
To this end, we have the following theorem.
Theorem 5.2.
Let $G$ be an $n$ node arbitrary, simple connected graph with a maximum degree $\Delta$ and diameter $D$. Let $n$ mobile agents with distinct IDs in the range $[0,n^{c}]$ with the highest ID $\lambda$, where $c$ is an arbitrary constant, be placed at each $n$ node of $G$ in an initial dispersed configuration. Then, the triangle centrality of each node $v\in G$ can be calculated in $O(\Delta\log\lambda)$ rounds if $\mathbf{T(G)}$ is known and in $O(D\Delta\log\lambda)$ rounds, if $\mathbf{T}(G)$ is unknown. $\mathbf{T}(G)$ is the total triangle count of the graph $G$.
5.3 Local Clustering Coefficient
The local clustering coefficient of a node in a graph is used to quantify how close its neighbours are to being a clique (complete graph) i.e., how well connected the network is around a particular node. Mathematically, the local clustering coefficient ($LCC$) of a node $v\in G$ can be written as $LCC(v)=\frac{T(v)}{\delta(v)(\delta(v)-1)}$,(from  [45]). Here $\delta(v)$ denotes the degree of the node $v$. This metric can be easily calculated using mobile agents. $T(v)$ for a node $v$ can be calculated in $O(\Delta\log\lambda)$ rounds [Lemma 3]. The agent which already knows $\delta(v)$ at $v$ can now evaluate $LCC(v)$ using the formula.
Theorem 5.3.
Let $G$ be an $n$ node arbitrary, simple connected graph with a maximum degree $\Delta$ and diameter $D$. Let $n$ mobile agents with distinct IDs in the range $[0,n^{c}]$ with the highest ID $\lambda$, where $c$ is an arbitrary constant, be placed at each $n$ node of $G$ in an initial dispersed configuration. Then, the Local Clustering Coefficient of each node $v\in G$, $LCC(v)$ can be calculated in $O(\Delta\log\lambda)$ rounds.
6 Conclusion and Future Work
In this paper, we introduced algorithms for counting triangles via mobile agents. The agents enumerated the triangles based on nodes and edges. Finally, the total number of triangles in the graph was calculated. We used the triangle enumeration algorithm and applied it to the Truss Decomposition Problem, and to calculate the Triangle Centrality and Local Clustering Coefficient for each node.
It would interesting to establish lower bounds for the above computations in terms of both time complexity and memory per robot. Also, a lower bound on the number of robots used to execute the above algorithm can be an interesting investigation. Furthermore, the efficiency of the problem in the faulty robot model can be another task to consider in future.

### v2

@@
-
+Prabhat Kumar Chand
 Abstract
-Triangle counting in a graph is a fundamental problem and has a wide range of applications in various domains.
-It is crucial in understanding the structural properties of a graph and is often used as a building block for more complex graph analytics.
-In this paper, we solve the triangle counting problem in an anonymous graph in a distributed setting using mobile agents and subsequently use this as a subroutine to tackle the truss decomposition and triangle centrality problem.
-The paper employs mobile agents, placed on the nodes of the graph to coordinate among themselves to solve the triangle enumeration problem for the graph.
-Following the literature, we consider the synchronous systems where each robot executes its tasks concurrently with all others and hence time complexity can be measured as the number of rounds needed to complete the task.
-The graph is anonymous, i.e., without any node labels or IDs, but the agents are autonomous with distinct IDs and have limited memory.
-Agents can only communicate with other agents locally i.e., if and only if they are at the same node.
-The goal is to devise algorithms that minimise both the time required for triangle counting and the memory usage at each agent.
-We further demonstrate how the triangle count obtained through the mobile agent approach can be leveraged to address the truss decomposition, triangle centrality and local clustering coefficient problems, which involves finding maximal sub-graphs with strong interconnections.
-Truss decomposition helps in identifying maximal, highly interconnected sub-graphs, or trusses, within a network, thus, revealing the structural cohesion and tight-knit communities in complex graphs, facilitating the analysis of relationships and information flow in various fields, such as social networks, biology, and recommendation systems.
-Keywords:
-Mobile Agents Triangle Counting $k$-Truss Truss Decomposition Triangle Centrality Local Clustering Coefficient Time Complexity Memory Complexity Network Algorithms Distributed Algorithms
+Triangle counting in a graph is a fundamental problem with wide-ranging applications across various domains.
+It is crucial for understanding graph structure and serves as a basis for more advanced graph analytics.
+One key application of triangle counting is truss decomposition, a technique that helps in identifying maximal, highly interconnected sub-graphs, or trusses, within a network, revealing the structural cohesion and tight-knit communities in complex graphs, facilitating the analysis of relationships and information flow in various fields, such as social networks, biology, and recommendation systems.
+Utilizing mobile agents or robots for tasks like truss decomposition and clustering coefficient calculations can be particularly advantageous in decentralized environments with limited, unreliable, or compromised communication systems.
+In such scenarios, these agents can carry out local computations, bypassing the need for extensive communication channels.
+This approach proves valuable in contexts such as disaster response, urban management, and military operations, where broadcast communication may be impractical.
+In this paper, we first address the triangle counting problem in an arbitrary anonymous graph using mobile agents.
+This method is then extended as a subroutine to solve the truss decomposition problem and compute the triangle centrality metric and local clustering coefficient for each node.
+Our approach utilizes $n$ autonomous mobile agents, each initially positioned at a different node of the $n$ node graph.
+These agents coordinate to collaboratively first solve the triangle enumeration problem and then truss decomposition and compute the triangle centrality and local clustering coefficient metrics.
+In line with existing research, we assume a synchronous system where all agents execute their tasks concurrently, allowing time complexity to be measured in rounds.
+The graph is anonymous, where the nodes lack any labels or IDs.
+However, the agents possess distinct IDs and some limited memory.
+Although the agents can perform extensive local computations, they can only communicate when they occupy the same node.
+Our primary goal is to design algorithms that minimize both the time required (which also reflects the number of agent movements) for triangle counting and the memory usage of each agent.
+Furthermore, we demonstrate how the triangle counts obtained through the mobile agent approach can be leveraged to solve the truss decomposition, triangle centrality, and local clustering coefficient problems.
+Keywords Agent-Based Systems, Autonomous Agents, Distributed Robot Systems, Multi-Robot Systems, Mobile Agents, Triangle Counting, $k$-Truss, Truss Decomposition, Triangle Centrality, Local Clustering Coefficient, Time Complexity, Network Algorithms, Distributed Algorithms
 1 Introduction
 Counting and listing triangles in a graph has received much attention in the last couple of decades as it serves as a building block of complex network analysis [2, 3].
-The number of triangles in a graph is used for computing the clustering coefficient, one of the most used metrics for network analysis [54, 7], and triangle centrality [37, 1, 8].
+The triangle count in a graph is used for computing clustering coefficient, one of the most used metrics for network analysis [2, 3], and triangle centrality [4, 5, 6].
 Triangle counting also plays a pivotal role in the hierarchical decomposition of a graph such as truss decomposition [7] which is an important hierarchical subgraph structure in community detection [8, 9].
-Triangle counting is used in solving many practical applications.
+Triangle counting is useful in many practical applications.
 Becchetti et al.
@@
 Other applications of triangle counting include query optimization in databases [12], link prediction in social network [13], and community detection in system biology [14].
-In this paper, we are interested in the triangle counting problem along with its applications in truss decomposition, computing triangle centrality and local clustering coefficient using autonomous agents (or robots) on anonymous graphs.
+In this paper, we are interested in the triangle counting problem along with its applications in truss decomposition, computing triangle centrality and local clustering coefficient using autonomous agents on anonymous graphs.
 Suppose we are given $n$ agents positioned initially on the nodes (each node has one agent) of an $n$-node anonymous graph $G$.
-The agents coordinate among themselves to solve the triangle counting problem such that each agent (at a node) outputs (i) node-based triangle count, (ii) edge-based triangle count and the total number of triangles in the graph (see the problem statements in Section 3).
+The agents coordinate among themselves to solve the triangle counting problem such that each agent (at a node) outputs (i) a node-based triangle count, (ii) an edge-based triangle count, and the total number of triangles in the graph (see the problem statements in Section 3).
 Our agent-based model has been gaining significant attention recently.
-For example, there have been some recent works on how to position the agents on nodes of the graph $G$ such that each agent’s position collectively form the maximal independent set (MIS) of $G$ [44, 42] or they identify a small dominating set [11] of $G$.
-Another related problem is of dispersion in which $k\leq n$ agents are positioned on $k$ different nodes of $G$, see [34] and the references therein.
-A solution to the dispersion problem guarantees that $k$ agents are positioned on $k$ different nodes; which is a requirement for the triangle counting problem defined in this paper.
-Exploration problem on graphs using mobile agents refers to solving a graph analytic task using one or more agents [16].
+For example, there have been some recent work on how to position the agents on nodes of the graph $G$ such that each agent’s position collectively forms the maximal independent set (MIS) of $G$ [15, 16] or they identify a small dominating set [17] of $G$.
+Recently, a similar mobile agent model has been used to explore the complexity of constructing a breadth-first search tree [18] and a minimum spanning tree in arbitrary anonymous graphs [19].
+Another related problem is of dispersion in which $k\leq n$ agents have to position themselves at $k$ different nodes of $G$, see [20] and the references therein.
+A solution to the dispersion problem guarantees that $k$ agents are positioned on $k$ different nodes; this is a requirement for the triangle counting problem defined in this paper.
+Exploration problems on graphs using mobile agents refer to solving a graph analytic task using one or more agents [21].
 In this work, we consider triangle counting in a simple, undirected, anonymous graph using mobile agents.
-The motivation stems from scenarios like private networks in the military or sensor networks in inaccessible terrain where direct access to the network is obstructed, but small battery-powered agents can navigate to learn network structures and their properties for overall network management.
-Prominent use of agents in network exploration can be seen in areas such as underwater navigation [15], network-centric warfare in military systems [36], modelling social network [56], studying social epidemiology [19] etc.
+The motivation comes from scenarios such as private networks in the military or sensor networks in inaccessible terrain where direct access to the network is obstructed, but small battery-powered agents can navigate to learn network structures and their properties for overall network management.
+The prominent use of agents in network exploration can be seen in areas such as underwater navigation [22], network-centric warfare in military systems [23], modeling social networks [24], studying social epidemiology [25] etc.
+In addition, triangle counting can be effectively applied to drone-based Social Network Analysis (SNA) in scenarios such as drone-assisted event monitoring at public gatherings, protests, or concerts.
+For example, as demonstrated in [26], drones can identify individuals in close proximity, such as those not wearing masks, and issue social distancing alerts through speakers.
+In addition, drones can collect data on interactions between individuals and triangle counting can be used to identify closely knit social groups, discover group dynamics, detect leaders, or monitor potentially suspicious clusters in real time, enhancing situational awareness and crowd management.
+Furthermore, mobile agents, such as drones, can play a crucial role in Search and Rescue (SAR) operations [27, 28, 29, 30].
+In SAR missions, multiple drones can cover large areas, locate missing persons, and maintain communication with the command centre.
+Drones can form resilient triangular communication networks where triangle counting helps identify strong communication links.
+This ensures robust communication paths between drones and the base station, facilitating quick information sharing.
+Drones are also being used for Wireless Sensor Network (WSN) deployment in remote or difficult-to-access areas, especially for environmental monitoring [31, 32, 33, 34].
+By identifying dense clusters of sensors or drones, triangle counting enhances network robustness.
+It helps detect areas of high data redundancy, optimize sensor placement, and improve communication resilience.
 1.1 Our Contributions
-We first enumerate the triangles in the graph $G$ and apply the triangle counting methodology to $3$ applications: (i) Truss Decomposition, (ii) Triangle Centrality, and (iii) Local Clustering Coefficient.
-Let $G$ be an $n$-node undirected and connected graph with the maximum degree $\Delta$ and diameter $D$.
-Let $n$ mobile agents with distinct IDs in the range $[0,n^{c}]$ ($c$ being an arbitrary constant) with the highest agent ID $\lambda\in[0,n^{c}]$, be placed at each of the $n$ nodes of $G$ in a dispersed initial configuration.
+We first enumerate the triangles in the graph $G$ and apply the triangle counting methodology to three applications: (i) Truss Decomposition, (ii) Triangle Centrality, and (iii) Local Clustering Coefficient.
+Let $G$ be an $n$-node simple, undirected, anonymous and connected graph with the maximum degree $\Delta$ and diameter $D$.
+Let $n$ mobile agents with distinct IDs with the highest agent ID $\lambda$, be placed at each of the $n$ nodes of $G$ in a dispersed initial configuration.
 Then, we solve the following problems.
@@
     • $\bullet$
-      Each agent $r_{i}$ can calculate and output the number of triangles adjacent to the vertex where $r_{i}$ is placed on, in $O(\Delta\log\lambda)$ rounds.
+      Each agent $r_{i}$ can calculate and output the number of triangles containing the vertex where $r_{i}$ is placed on, in $O(\Delta\log\lambda)$ rounds.
     • $\bullet$
-      Each agent $r_{i}$ can calculate the number of triangles based on each of its adjacent edges in $O(\Delta\log\lambda)$ rounds.
+      Each agent $r_{i}$ can calculate the number of triangles based on each edge of $G$ in $O(\Delta\log\lambda)$ rounds.
     • $\bullet$
-      Each agent $r_{i}$ can calculate the number of triangles in $G$ in $O(D\Delta\log\lambda)$ rounds.
+      Each agent $r_{i}$ can calculate the total number of triangles in $G$ in $O(D\Delta\log\lambda)$ rounds.
     The memory requirement for each agent is $O(\Delta\log n)$ bits.
   2. 2.
-    Truss Decomposition - The Truss Decomposition Problem for $G$ can be solved by the mobile agents in $O(m\Delta D\log\lambda)$ rounds with $O(\Delta\log n)$ bits of memory per agent.
-  3. 3.
+    Applications
+    • $\bullet$
+      Truss Decomposition - The Truss Decomposition Problem for $G$ can be solved by the mobile agents in $O(m\Delta\log\lambda+mD)$ rounds.
+    • $\bullet$
       Triangle Centrality - The Triangle Centrality of each node $v\in G$ can be calculated in $O(\Delta\log\lambda)$ rounds if $\mathbf{T(G)}$ is known and in $O(D\Delta\log\lambda)$ rounds, if $\mathbf{T}(G)$ is unknown.
 $\mathbf{T}(G)$ is the total triangle count of the graph $G$.
-  4. 4.
+    • $\bullet$
       Local Clustering Coefficient - The Local Clustering Coefficient of each node $v\in G$, i.e., $LCC(v)$ can be calculated by the $n$ agents in $O(\Delta\log\lambda)$ rounds.
+    The memory requirement for each agent is $O(\Delta\log n)$ bits.
 The notations are from Table 1.
-1.2 Possible Applications
-Triangle Counting and Truss Decomposition, in general, have various applications in community detection, social network analysis, biological networks, recommendation systems etc.
-Mobile agents on the other hand, equipped with truss decomposition capabilities, post dispersion, can dynamically identify and navigate through structurally significant regions of the graph.
-For example, an unmanned mobile agent can help prioritize areas within a large area that require prompt rescue and can carry rescue materials to difficult-to-access terrain.
-Similarly, robots carrying different agricultural materials (like fertilizers or pesticides) can dynamically arrange themselves and focus on specific areas that exhibit particular structural characteristics relevant to crop health.
-Truss decomposition can contribute to the autonomy of vehicles by helping them navigate efficiently through road networks or complex urban environments or helping them identify core areas in a city.
-In general, mobile robots can prioritize exploring areas with high truss density, indicating regions of potential interest or connectivity within the environment.
-In addition, its ability to modify itself in challenging scenarios improves fault tolerance and helps maintain the connectivity of the network.
 2 Related Work
@@
 2.1.1 Sequential Algorithms:
-In 1985, Chiba and Nishizeki [13] proposed an algorithm for counting all triangles in a simple graph by computing the intersection of the neighborhoods of the adjacent vertices with time complexity $O(m^{\frac{3}{2}})$ where $m$ is the number of edges in the graph.
+In 1985, Chiba and Nishizeki [35] proposed an algorithm for counting all triangles in a simple graph by computing the intersection of the neighbourhoods of the adjacent vertices with time complexity $O(m^{\frac{3}{2}})$ where $m$ is the number of edges in the graph.
 Other sequential algorithms for triangle counting are based on vertex-iterator [36] and edge-iterator [37].
-In vertex iterator based technique, it iterates over each vertex $v$ of the graph and intersects the adjacency list of each pair of the neighbors of $v$.
-In edge-iterator based technique, it iterates over each edge and intersects the adjacency list of its two endpoints.
+In the vertex iterator based technique, it iterates over each vertex $v$ of the graph and intersects the adjacency list of each pair of the neighbours of $v$.
+In the edge-iterator based technique, it iterates over each edge and intersects the adjacency list of its two endpoints.
 2.1.2 Parallel Algorithms:
@@
 2.2 Truss Decomposition
-In [14], Cohen introduced and visualized truss as a relaxation to cliques in graphs and defined it to be a non-trivial, one-component sub-graph such that each edge is reinforced by at least $k-2$ pairs (for a $k$-truss) of edges making a triangle with that edge.
+In [48], Cohen introduced and visualized truss as a relaxation to cliques in graphs and defined it to be a non-trivial, one-component sub-graph such that each edge is reinforced by at least $k-2$ pairs (for a $k$-truss) of edges, making a triangle with that edge.
 He established the existence of polynomial time algorithms for identifying $k$-trusses within a given graph.
-Since then, trusses have become the subject of extensive investigation the problem has been studied in various models  [24, 12, 25, 29].
+Since then, trusses have become the subject of extensive investigation; the problem has been studied in various models  [8, 49, 50, 51].
 In general, truss decomposition is explored in two main contexts: a serialized version, primarily suitable for small to medium-sized graphs, and a parallel version designed for handling larger graphs.
@@
 In particular, their algorithm computed the $k-$trusses for all $k\geq 3$ in $O(m^{1.5})$ time using $O(m+n)$ memory space.
-They further proposed two I/O efficient algorithms for handling massive networks which are ill-equipped for the main memory of a single machine.
+They further proposed two I/O efficient algorithms for handling massive networks, which are ill-equipped for the main memory of a single machine.
 To tackle the large memory requirement for massive graphs, several parallel implementations for truss decomposition were introduced.
@@
 They developed parallel algorithms for both synchronous and asynchronous versions of the nucleus decomposition problem.
-In the synchronous version, the computations for specific iterations are made over a single snapshot of current values whereas in the asynchronous version, the latest values of each variable is used instantly in the current computation.
+In the synchronous version, the computations for specific iterations are made over a single snapshot of current values, whereas in the asynchronous version, the latest values of each variable is used instantly in the current computation.
 Voegele et al.
@@
 The efficiency of most of the algorithms is described with the aid of extensive experimental results.
-Agent-based Computations on Graphs: In [49], Sudo et al.
+2.3 Agent-based Computations on Graphs:
+In [57], Sudo et al.
 considered the exploration problem with a single agent in undirected graphs.
 Starting from an arbitrary node, the agent has to explore all the nodes and edges in the graph and return to the starting node.
-The authors used a whiteboard model, that reduced the memory requirement per robot.
-In  [30], the authors explored the graph exploration problem based on Depth-First-Search and studied the trade-off between node-memory vs robot-memory.
+The authors used a whiteboard model, that reduced the memory requirement per agent.
+In  [58], the authors explored the graph exploration problem based on Depth-First-Search and studied the trade-off between node-memory vs.
+agent-memory.
 In [59], Dereniowski et al.
@@
 They also obtained almost tight bounds on the asymptotic relation between exploration time and team size, for large $k$, in both the local and the global communication model.
-Further results on agent-based graph exploration and dispersion (spreading agents across the graph so that there are at-most one agent at each node) can be found in several recent papers [31, 32, 33, 39, 40, 43, 10].
+Further results on agent-based graph exploration and dispersion (spreading agents across the graph so that there is at-most one agent at each node) can be found in several recent papers [60, 61, 62, 63, 64, 65, 66].
 3 Model and Problem Definitions
@@
 Any number of agents can travel through an edge at any given time.
-Mobile Agents (or Mobile Robots): We have a collection of $n$ agents =$\mathcal{R}=\{r_{1},r_{2},...,r_{n}\}$ residing on the nodes of the graph in such a way that each node is occupied by a distinct ID agent at the start (known as dispersed configuration in literature).
-Each agent has a unique ID in the range $[0,n^{c}]$ ($c$ is an arbitrary constant) and has $O(\Delta\cdot\log(n))$ bits to store information.
-An agent therefore cannot store the whole graph structure information within its limited memory.
-An agent retains its memory as long as needed and it can be updated as required.
+Mobile Agents: We have a collection of $n$ agents =$\mathcal{R}=\{r_{1},r_{2},...,r_{n}\}$ residing on the nodes of the graph in such a way that each node is occupied by a distinct ID agent at the start (known as dispersed configuration in literature).
+Each agent has a unique ID (we bound the IDs of the agents by $n^{c}$, for some large constant $c$, to accommodate the ID space of the agents within an $O(\log n)$ bit ID field) and has a limited memory to store information.
+An agent cannot store the whole graph structure information within its memory.
+An agent retains its memory as long as needed, and it can be updated as required.
 Two or more agents can be present (co-located) at a node or pass through an edge in $G$.
-However, an agent is not allowed to stay on an edge.
+However, an agent is not allowed to stay on edge.
 An agent can recognise the port number through which it has entered and exited a node.
@@
 Only the collocated agents at a node can sense each other and exchange information.
+Every agent has prior knowledge of the highest degree of the graph $\Delta$ and its diameter $D$.
+The agents are also aware of the highest ID among themselves, denoted as $\lambda$.
 An agent can transfer all the information stored in its memory in a single round.
-We interchangeably use the terms robots and agents.
 Communication Model: We consider a synchronous system where the agents are synchronised to a common clock.
@@
 $n,m$ Number of nodes and edges of G respectively
-$\mathcal{R}$ Collection of $n$ mobile robots $\{r_{1},r_{2},\dots,r_{n}\}$
-$\lambda$ ID of the robot with maximum ID
+$\mathcal{R}$ Collection of $n$ mobile agents $\{r_{1},r_{2},\dots,r_{n}\}$
+$\lambda$ ID of the agent with maximum ID
 $\delta(v)$ Degree of the node $v$
@@
 3.1 Problem Statements
-Problem 1.
-(Triangle Counting using Mobile Agents) Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ agents, each of which initially placed distinctly at each node of $G$.
+Triangle Counting using Mobile Agents: Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of $n$ agents, each of which initially placed distinctly at each node of $G$.
 We solve the following problems.
-  1. 1.
+  1. (a)
     Node-Based Triangle Counting: To count the number of triangles with a given node as a vertex.
-  2. 2.
+  2. (b)
     Edge-Based Triangle Counting: To count the number of triangles based on a given edge.
-  3. 3.
+  3. (c)
     Total Triangle Counting: To count the total number of triangles in the graph $G$.
-Problem 2.
-(Truss Decomposition using Mobile Agents) Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of $n$ agents, each of which are initially placed distinctly at each node of $G$.
-The $n$ autonomous agents coordinate among themselves to solve the Truss Decomposition Problem.
-Problem 3.
-(Triangle Centrality using Mobile Agents) Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of $n$ agents, each of which initially placed distinctly at each node of $G$.
-The $n$ autonomous agents coordinate among themselves to compute the Triangle Centrality $TC(v)$ for each node $v\in G$.
-Problem 4.
-(Local Clustering Coefficient using Mobile Agents) Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of $n$ agents, each of which initially placed distinctly at each node of $G$.
-The $n$ autonomous agents coordinate among themselves to compute the Local Clustering Coefficient, $LLC(v)$ for each node $v\in G$.
+Truss Decomposition, Triangle Centrality and Local Clustering Coefficient: Consider an undirected, simple, connected anonymous $n$-node graph $G=(V,E)$ and a collection =$\mathcal{R}=\{r_{1},r_{2},\dots,r_{n}\}$ of $n$ agents, each of which initially placed distinctly at each node of $G$.
+The $n$ autonomous agents coordinate among themselves to solve the (i) Truss Decomposition Problem and compute (ii) Triangle Centrality, and (iii) Local Clustering Coefficient of a given node.
 In this paper, we study the above problems from a theoretical perspective and aim to solve them while minimizing both time and memory-per-agent as much as possible.
@@
 In the first phase, the agents learn their neighbours.
-In the second phase, the agents check the number of common neighbours with each of its adjacent agents.
+In the second phase, the agents check the number of common neighbours with each of their adjacent agents.
 In this phase, each agent $r_{i}$ also counts the number of local triangles with $r_{i}$ as a vertex and the number of triangles with $(r_{i},r_{j})$ as an edge, where $r_{j}$ is an adjacent agent to $r_{i}$.
 In the third and last phase, each agent collects the local triangle count from every other agent and counts the number of triangles in $G$.
-We now explain each phase in detail, below.
-4.1 Phase 1: Know the Neighbourhood
-Each node is occupied by a distinct (via their IDs) agent before the start of the algorithm and we represent each node of $G$ by its stationed agent.
+We now explain each phase in detail below.
+4.1 Phase 1: Neighbourhood Discovery (Protocol MYN)
+Each node is occupied by a distinct (via their IDs) agent before the start of the algorithm, and we represent each node of $G$ by its stationed agent.
 The algorithm starts with Phase 1, where each agent discovers and records its neighbour.
@@
 Therefore, we need to ensure that each gets to record the correct neighbour set of it.
-To do this we exploit the ID bits of the agents.
+To do this, we exploit the ID bits of the agents.
 We use the fact the IDs of the agents are distinct.
 To this end, we state the following lemma.
-Lemma 1.
-Let $r_{i}$ and $r_{j}$ be two distinct agents in $\mathcal{R}$, with their IDs field being $r_{i}.ID$ and $r_{j}.ID$ respectively.
+Lemma 4.1.
+Let $r_{i}$ and $r_{j}$ be two distinct agents in $\mathcal{R}$, with their IDs fields being $r_{i}.ID$ and $r_{j}.ID$ respectively.
 Then, there exists at least one dissimilar bit in $r_{i}.ID$ and $r_{j}.ID$ with one being 00 and the other being $1$.
@@
 After the completion of Phase 1 of our algorithm, it is guaranteed that each agent correctly records all its neighbouring agents in $O(\Delta\cdot\log\lambda)$ rounds, which we prove in the following lemma.
-Lemma 2.
+Lemma 4.2.
 Each agent $r_{i}$ correctly records the exhaustive list of its neighbour agents in $O(\Delta\cdot\log\lambda)$ rounds after the completion of Phase 1.
@@
 Let $r_{j}$ (placed at node $v$) be a neighbour of $r_{i}$ (at node $u$).
-Then there exists a port joining $u$ to $v$ having port number between $[0,\delta(u)-1]$.
-Since $r_{i}$ and $r_{j}$ have distinct IDs, there exists a bit (say the $p^{th}$ bit from the right) in $r_{i}.ID$ and $r_{j}.ID$, which are different from one another (one bit being 00, the other being $1$)[Lemma 1].
+Then there exists a port joining $u$ to $v$ having a port number between $[0,\delta(u)-1]$.
+Since $r_{i}$ and $r_{j}$ have distinct IDs, there exists a bit (say the $p^{th}$ bit from the right) in $r_{i}.ID$ and $r_{j}.ID$, which are different from one another (one bit being 00, the other being $1$)[Lemma 4.1].
 Without the loss of generality, let us assume that the $p^{th}$ bit of $r_{i}$ is $1$ and the $p^{th}$ bit of $r_{i}$ is 00.
@@
 Therefore, with the end of Phase 1, each agent has now enlisted its neighbouring agents in its memory.
+We call this particular step Protocol MYN, a protocol for Meeting-Your-Neighbour.
 The algorithm now moves to Phase 2 where each agent $r_{i}$ counts the number of triangles with $r_{i}$ as one of its vertex (node).
 4.2 Phase 2: Local Triangle Counting
-In phase two, each agent, now equipped with the list of its neighbouring agents, visits each of its neighbours once again, in exactly the similar fashion described in Phase 1.
+In Phase 2, each agent, now equipped with the list of its neighbouring agents, visits each of its neighbours once again, in exactly the same manner as described in Phase 1.
 Whenever the agent $r_{i}$ meets its neighbour $r_{j}$, they communicate to find out the number of common neighbours they both have.
-Here as the nodes are anonymous, they are identified using the mobile agents that reside on the nodes.
-This communication is also used the update the following variables of $r_{i}$:
+Since the nodes are anonymous, they are identified using the mobile agents that reside on the nodes.
+This communication is also used to update the following variables of $r_{i}$:
   1. 1.
@@
 $r_{i}$ has $\Delta$ such variables $r_{i}.edge(x)$, where $x$ represents another agent that has an edge with $r_{i}$.
-Each of these variables are initially set to 00.
+This, therefore requires an additional $O(\Delta\cdot\log n)$ memory per agent which forms the dominant factor in the memory requirement.
+Each of these variables is initially set to 00.
 A variable $r_{i}.edge(x)$ remains 00, if there are no triangles with $(r_{i},x)$ as an edge.
   2. 2.
-    $r_{i}.local\_sum:$ Initially set to 00, adds up the counts of the number of common neighbours for each distinct neighbour of $r_{i}$.
+    $r_{i}.local\_sum:$ Initially set to 00, adds up the counts of the common neighbours for each distinct neighbour of $r_{i}$.
 As $r_{i}$ finds its neighbours one by one, it cumulatively adds up the count of the number of common neighbours with each of its neighbours.
@@
 We state the reason for the same in the following lemma.
-Lemma 3.
+Lemma 4.3.
 The number of triangles with agent $r_{i}$ as a vertex is given by $\frac{1}{2}(r_{i}.local\_sum)$.
 Proof.
-First, we see that if the agents $r_{i}$ and $r_{j}$ have a common neighbour $r_{k}$, then, $(r_{i},r_{j},r_{k})$ form a triangle with $r_{i}$ as one of the vertex.
+First, we observe that if the agents $r_{i}$ and $r_{j}$ have a common neighbour $r_{k}$, then, $(r_{i},r_{j},r_{k})$ form a triangle with $r_{i}$ as one of the vertices.
 Now, the triangle $(r_{i},r_{j},r_{k})$ has been tallied once while counting the common neighbour of $r_{i}$ with $r_{j}$ and re-counted again as a common neighbour of $r_{i}$ with $r_{k}$.
-Therefore, each $(r_{i},r_{j},r_{k})$ with $r_{i}$ as vertex, is counted twice, once through $r_{j}$ and other time through $r_{k}$.
+Therefore, each $(r_{i},r_{j},r_{k})$ with $r_{i}$ as vertex, is counted twice, once through $r_{j}$ and the other time through $r_{k}$.
 So, the number of triangles with agent $r_{i}$ as a vertex is exactly half of $r_{i}.local\_sum$.
 ∎
-Therefore at the end of Phase 2, each agent $r_{i}$ gets a local count of the number of triangles with $r_{i}$ as a vertex and a list of the number of triangles that are based on the edges adjacent to $r_{i}$.
-In the next phase, the algorithm finds the number of triangles of $G$ using the counts generated in this phase.
+Thus, at the end of Phase 2, each agent $r_{i}$ gets a local count of the number of triangles with $r_{i}$ as a vertex and a list of the number of triangles that are based on the edges adjacent to $r_{i}$.
+In the next phase, the algorithm determines the number of triangles of $G$ using the counts generated in this phase.
 4.3 Phase 3: Counting the Number of Triangles in $G$
 To find the number of triangles in $G$, we need to take into account the local triangle counts of each agent.
-We first need to accumulate the $r_{i}.local\_sum$ values of each of the $n$ agents and calculate the triangles of $G$ from there.
+First, we need to accumulate the $r_{i}.local\_sum$ values of each of the $n$ agents and calculate the triangles of $G$ from there.
 However, due to the absence of a $leader$ agent and the difficulty of synchronising the movement of the agents, gathering the $r_{i}.local\_sum$ is not straightforward.
-The high-level idea is to communicate repeatedly with the neighbours and continually gather up the values of $local\_sum$ of each agent hop by hop till each agent gets every $local\_sum$ value.
-As described in the previous case, the agents can assuredly communicate with all their neighbours in $\Delta\cdot\log\lambda$ rounds.
-Inside the first $\Delta\cdot\log\lambda$ round of this phase, the agents communicate with their neighbours and exchange the variable $local\_sum$.
-An agent $r_{i}$, along with its own $r_{i}.local\_sum$, collects $r_{j}.local\_sum$ values for each of its neighbour $r_{j}$ and stores it in the memory.
-Within the next $\Delta\cdot\log\lambda$ rounds, $r_{i}$ meets its neighbour agents $r_{j}$ again to check if they have collected any new $local\_sum$ values (possibly from their own neighbour).
-$r_{i}$ again stores any new $local\_sum$ record (of its neighbour’s neighbour) that it receives through its neighbour $r_{j}$ in the current phase.
+The high-level idea is to communicate repeatedly with the neighbours and continually gather up the values of $local\_sum$ of each agent, hop by hop, until each agent gets every $local\_sum$ value.
+As described previously, the agents can reliably communicate with all their neighbours in $\Delta\cdot\log\lambda$ rounds.
+During the first $\Delta\cdot\log\lambda$ round of this phase, the agents communicate with their neighbours and exchange their $local\_sum$ values.
+An agent $r_{i}$, along with its own $r_{i}.local\_sum$, collects $r_{j}.local\_sum$ values from each of its neighbours $r_{j}$ and stores them in memory.
+During the next $\Delta\cdot\log\lambda$ rounds, $r_{i}$ meets its neighbour agents $r_{j}$ again to check if they have collected any new $local\_sum$ values (possibly from their own neighbour).
+$r_{i}$ again stores any new $local\_sum$ record (of its neighbour’s neighbour) that it receives through its neighbour $r_{j}$ during this phase.
 By the end of the second phase, $r_{i}$ has the information of the $local\_sum$ of all the agents that are at a distance of $2-hops$ from it.
-Now, with the completion of $D$ (the diameter of the graph $G$) such $\Delta\cdot\log\lambda$ rounds, each agent $r_{i}$ now has the $local\_sum$ record of every agent in $G$.
+Now, with the completion of $D$ (the diameter of the graph $G$, which is known to the agents) such $\Delta\cdot\log\lambda$ rounds, each agent $r_{i}$ now has the $local\_sum$ record of every agent in $G$.
 The number of triangles in $G$ is finally calculated by each agent by summing up the $n$ $local\_sum$ values that they have gathered and dividing it by $6$.
-Lemma 4.
+Lemma 4.4.
 Let $r_{i}$ and $r_{k}$ be two agents at a distance of $d$ hops from each other. Then, the value of $r_{k}.local\_sum$ can be communicated to $r_{j}$ within $d\cdot O(\Delta\cdot\log\lambda)$ rounds.
@@
 In the first $O(\Delta\cdot\log\lambda)$ rounds, $r_{i}$ can communicate with $r_{i_{1}}$ to get the value of $r_{i_{1}}.local\_sum$.
-In the meanwhile, $r_{i_{1}}$ also collects the value of $r_{i_{2}}.local\_sum$ in the same $O(\Delta\cdot\log\lambda)$ round.
+Meanwhile, $r_{i_{1}}$ also collects the value of $r_{i_{2}}.local\_sum$ during the same $O(\Delta\cdot\log\lambda)$ round.
 Therefore, in the second sub-phase of $O(\Delta\cdot\log\lambda)$ rounds, when $r_{i}$ communicates with $r_{i_{1}}$ again, $r_{i}$ receives the value of $r_{i_{2}}.local\_sum$ through $r_{i_{1}}$.
@@
 ∎
-Let $D$ denote the diameter of the graph $G$. Since any two agents in $G$ are located at a distance of at most $D$ hops, the following lemma follows from Lemma 4.
-Lemma 5.
-After the end of Phase 3, which takes $O(D\cdot\Delta\cdot\log\lambda)$ rounds, each agent $r_{i}$ has the complete record of $local\_sum$ values of every agent in $G$.
-Here, $D$ is the diameter of the graph $G$.
-If the diameter $D$ of the graph is unknown, $D$ could be replaced by $n$.
-Lemma 6.
+Let $D$ denote the diameter of the graph $G$. Since any two agents in $G$ are at-most $D$ hops apart, the following lemma follows from Lemma 4.4.
+Lemma 4.5.
+By the end of Phase 3, which takes $O(D\cdot\Delta\cdot\log\lambda)$ rounds, each agent $r_{i}$ has the complete record of $local\_sum$ values of every agent in $G$.
+Lemma 4.6.
 The number of triangles in graph $G$ is given by $\frac{1}{6}\cdot\Sigma_{i}(r_{i}.local\_sum)$, where the sum runs over all the $n$ agents of $G$.
@@
 Notation: We shall denote the number of triangle containing the vertex (node) $v$ and the total number of triangles in $G$ with $\mathbf{T}(v)$ and $\mathbf{T}(G)$, respectively.
-For a node $v$ with a robot $r_{i}$ on it, we use $r_{i}$ and $v$ interchangeably to denote the node.
-At the end of this phase, each agent has the value of the number of triangles in $G$ i.e., $\mathbf{T}(G)$. In the following theorem, we assemble the list of results we got during the 3 phases.
+For a node $v$ with an agent $r_{i}$ on it, we use $r_{i}$ and $v$ interchangeably to denote the node.
+At the end of this phase, each agent has the value of the number of triangles in $G$ i.e., $\mathbf{T}(G)$. In the following theorem, we assemble the list of results we obtained during the 3 phases.
 Theorem 4.7.
@@
   3. 3.
-    Each agent $r_{i}$ can calculate the number of triangles in $G$, $\mathbf{T}(G)$, in $O(D\cdot\Delta\cdot\log\lambda)$ rounds.
+    Each agent $r_{i}$ can calculate the number of triangles in $G$, $\mathbf{T}(G)$, in $O(D\cdot\Delta\cdot\log\lambda)$ rounds,
+using $O(\Delta\log n)$ bits of memory per agent.
 5 Applications: Truss Decomposition, Triangle Centrality and Local Clustering Coefficient
-In this section, we show some novel applications of our triangle-counting algorithm via mobile agents in an anonymous graph.
+In this section, we show some applications of our triangle-counting algorithm via mobile agents in an anonymous graph, namely in solving the truss decomposition problem and computing the triangle centrality metric and local clustering coefficient for each node of the graph.
+Truss decomposition, and clustering coefficient calculations using mobile agents offer significant advantages in environments where decentralized and adaptive operations are crucial.
+In scenarios with limited or unreliable communication infrastructure, mobile agents can perform local computations, reducing the need for extensive data transmission.
+This capability is especially valuable in temporary or ad hoc networks, such as those formed during disaster response, where agents can quickly analyse the network’s structure to identify robust sub-graphs or highly connected areas, ensuring that resources are efficiently allocated.
+For instance, drones operating in disaster zones can perform truss decomposition to pinpoint areas of strong connectivity, directing rescue efforts where the network of survivors or critical infrastructure is most resilient.
+In urban environments, mobile agents representing vehicles can compute clustering coefficients to identify highly interconnected regions of a city, enabling more effective traffic management and resource distribution.
+In the military context, UAVs can employ triangle counting to reveal strategic communication hubs, improving surveillance and ensuring critical areas are secured against potential disruptions.
 5.1 Truss Decomposition
@@
 A sub-graph $T_{k}$ of $G$ is called a $k$-truss if every edge of $T_{k}$ is a part of at least $k-2$ triangles i.e, each edge in $T_{k}\subset G$ is supported by at least $k-2$ triangles in $T_{k}$.
-In our algorithm, we use the popular truss decomposition method to first calculate, what is called, $trussness$ value for each edge.
+In our algorithm, we use the popular truss decomposition method to first calculate what is called the $trussness$ value for each edge.
 The $trussness$ values are then used to construct the $k$-truss sub-graph for any value of $k$.
 Before we explain our algorithms in detail, we formally define $support$, $k$-truss and $trussness$.
-Definition 1 ($support$).
+Definition 5.1 ($support$).
 For a given graph $G(V,E)$, the $support$ of an edge $e\in E$ is the number of $triangles$ in $G$ that contain $e$.
-Definition 2 ($k$-truss).
+Definition 5.2 ($k$-truss).
 […]
-Definition 3 ($trussness$).
+Definition 5.3 ($trussness$).
 The $trussness$ of an edge $e$, is defined as the maximum $k$ such that $e$ belongs to $T_{k}$ but does not belong to $T_{k+1}$.
-Algorithm for Truss Decomposition: Let us now propose an algorithm for mobile agents to evaluate $trussness$ for each edge of the graph $G(V,E)$.
+Algorithm for Truss Decomposition: We now propose an algorithm for mobile agents to evaluate $trussness$ for each edge of the graph $G(V,E)$.
 The algorithm is based on the parallel truss decomposition described in [55].
-The $trussness$ values determine a partition (thus an equivalence relation) on $E$, where each class have the edges of $G$ with the same $trussness$ value.
+The $trussness$ values determine a partition (thus an equivalence relation) on $E$, where each class has the edges of $G$ with the same $trussness$ value.
 Let $\mathbf{t_{k}}$ denote the equivalence class of edges having $trussness=t_{k}$.
@@
 Mathematically, $T_{k}=\cup_{i=k}^{t_{max}}\mathbf{t_{i}}$.
-Therefore by computing the trussness for each edge in $G(V,E)$, we obtain a partition (equivalence classes) of $E$, thereby obtaining the $k$-trusses of $G$ for any $k$ by taking the union of the equivalent classes.
+Therefore, by computing the trussness for each edge in $G(V,E)$, we obtain a partition (equivalence classes) of $E$, thereby obtaining the $k$-trusses of $G$ for any $k$ by taking the union of the equivalent classes.
 Therefore, the $k$-truss decomposition of a graph is equivalent to computing the $trussness$ of each edge in the graph.
@@
 The removed edge (say $e$) keeps its final $support+2$ as its $trussness$ value.
-Once the $e$ is removed, the $support$ value of the edges that formed the triangle with $e$ is re-evaluated.
+Once $e$ is removed, the $support$ value of the edges that formed the triangle with $e$ is re-evaluated.
 The remaining edges are once again sorted in order before running another pass.
 The need to sort the edges in order of $support$ makes the algorithm inherently sequential.
-In the parallel version of the algorithm, the sorting condition is relaxed with the use of $h-index$ updating.
+In the parallel version of the algorithm, the sorting condition is relaxed with the use of $h-index$ upgradation.
 For a set of integers $S$, the $h-index$ of $S$ is defined as the largest number $h$ such that there are at least $h$ elements in $S$ that are equal to and greater than $h$.
@@
 The algorithm iteratively updates an edge’s $h-index$ by computing the $h-index$ of all edges that support it, until achieving convergence when no updates would happen.
-The final $h-index$ of each edge before, no further updates happen, provide the $trusness$ value for each edge.
+The final $h-index$ of each edge before no further updates happen, provide the $trusness$ value for each edge.
 However, mobile agents bring unique challenges.
@@
 $r_{i}$ stores only edges $(r_{i},r_{j})$, where $r_{j}>r_{i}$.
-$edge\_set$ requires $O(\Delta)$ bit of memory, where $\Delta$ is the highest degree of a node in $G$.
+$edge\_set$ requires $O(\Delta\log n)$ bit of memory, where $\Delta$ is the highest degree of a node in $G$.
 The ordering of edges ensures that each edge becomes associated with exactly one agent.
@@
   4. 4.
-    $L$ also associated with each edge in $edge\_set$.
-$L$ is used to store the values $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$ for for each triangle $(e,e^{\prime},e^{\prime\prime})$ containing $e$.
+    $L$ also associated with each edge in $edge\_set$. $L$ is used to store the values $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$ for each triangle $(e,e^{\prime},e^{\prime\prime})$ containing $e$.
   5. 5.
@@
   6. 6.
-    $scheduled$ to identify the edges in the $edge\_set$ that need further updating in consequent rounds of the algorithm. Initially, every edge in $edge\_set$ is scheduled ($true$).
+    $scheduled$ to identify the edges in the $edge\_set$ that need further updating in subsequent rounds of the algorithm. Initially, every edge in $edge\_set$ is scheduled ($true$).
   7. 7.
@@
 With each node of $G$ hosting a distinct agent, each agent first visits and registers its list of neighbours.
-Due to a lack of movement synchronisation, the agents follow the exact protocol as described in Phase 1 of Section 4 to register the neighbours, the ports which lead to the respective neighbours and the respective edges.
+Due to a lack of movement synchronisation, the agents follow the exact protocol as described in Phase 1 of Section 4 to register the neighbours, the ports that lead to the respective neighbours and the respective edges.
 To avoid inconsistency, the agents maintain the record of the edges in a specific order.
@@
 Let us assume, at some point of time during the algorithm, an agent $r_{i}$ meets its neighbour $r_{j}$.
-$r_{i}$ then checks the number of common neighbours with $r_{j}$ and records the count into $r_{i}.support(r_{i},r_{j})$ (we have assumed that $r_{j}.ID>r_{i}.ID$; otherwise if $r_{j}.ID<r_{i}.ID$, then the edge $(r_{i},r_{j})$ (simultaneously, $r_{i}.support(r_{i},r_{j})$) does not register in $r_{i}$’s records and in such case, the agent $r_{i}$ does nothing.
-However, in such case, the edge does register in $r_{j}$’s record, and $r_{j}$ then does the needful in that respect.).
+$r_{i}$ then checks the number of common neighbours with $r_{j}$ and records the count into $r_{i}.support(r_{i},r_{j})$ (we have assumed that $r_{j}.ID>r_{i}.ID$; otherwise if $r_{j}.ID<r_{i}.ID$, then the edge $(r_{i},r_{j})$ (simultaneously, $r_{i}.support(r_{i},r_{j})$) does not register in $r_{i}$’s records and in such a case, the agent $r_{i}$ does nothing.
+However, in such a case, the edge does register in $r_{j}$’s record, and $r_{j}$ then does the needful in that respect.).
 Support calculation takes $O(\Delta\log\lambda)$ rounds, after which, each agent receives its $support$ number.
-With the execution of this, $h$ value for each edge in $r_{i}.edge\_set$ gets initialized to the support of the particular edge.
+With the execution of this, the $h$ value for each edge in $r_{i}.edge\_set$ gets initialized to the support of the particular edge.
 5.1.3 Phase 3 (Iterative $h-index$ up-gradation)
@@
 The phase iterates till the $change$ value in any of the agents $r_{i}$ remains 00 and terminates once the value of $change$ in each and every agent becomes $1$.
-In each running phase, each agent $r_{i}$ does the following in parallel.
+In each phase, each agent $r_{i}$ does the following in parallel.
   1. 1.
     Checks if any edge is scheduled for further up-gradation in its $edge\_set$.
-If no, set $r_{i}.change\leftarrow 1$ and stops executing the algorithm but remains active.
+If no, set $r_{i}.change\leftarrow 1$ and stop executing the algorithm but remain active.
 $r_{i}$ provides any information when required to other agents.
@@
 Now for $r_{i}$ to calculate $\min\{h(e^{\prime}),h(e^{\prime\prime})\}$, $r_{i}$ must visit $r_{p}$.
-Since the movement of the agents is not centralised, $r_{i}$ must execute "Know the Neighbourhood" (Phase 1, Sec.
+Since the movement of the agents is not centralised, $r_{i}$ must execute ”Neighbourhood Discovery” (Phase 1, Sec.
 4.1) which takes additional $O(\Delta\log\lambda)$ rounds, in order to guarantee that $r_{i}$ meets $r_{p}$.
@@
 For each triangle $(e,e^{\prime},e^{\prime\prime})$ the agents add $e^{\prime},e^{\prime\prime}$ to their respective $N$ sets.
-Now, since each edge, $e^{\prime}$ or $e^{\prime\prime}$ is associated with a unique agent, an invocation of "Know the Neighbourhood" (Phase 1, Sec.
+Now, since each edge, $e^{\prime}$ or $e^{\prime\prime}$ is associated with a unique agent, an invocation of ”Protocol MYN” (Phase 1, Sec.
 4.1) is required to inform the correct agent associated with a particular edge so that the agent can schedule the particular edge in the next step if required.
@@
   6. 6.
-    Checking for Termination: Each agent, after having evaluated the $change$ value for the current phase, communicates with each of the other agents to check, if the algorithm is ready for termination.
-The agent sends the $change$ values to every agent in multiple rounds.
-To communicate, each agent starts by visiting its immediate neighbours.
-In the first round, each agent sends the $change$ to its neighbour.
-In the second round, the same change value is propagated to the 2-hop neighbour and so on.
-By the end of $D$ rounds, each agent receives every changed value from every other agent.
-Now, the agents calculate the binary product of all the $change$ values received.
-If the binary product of the $change$ value results in $1$ (all the agents must get the same binary product value), the agents terminate the algorithm.
-Otherwise, the agents rerun Phase 3 again, till the product becomes $1$.
-Lemma 7.
-Algorithm 5.1 takes $O(m\Delta D\log\lambda)$ rounds to terminate
+    Checking for Termination: To determine whether the algorithm is ready for termination, each agent first evaluates the $change$ value for the current phase.
+Each agent then communicates with all other agents to collect their $change$ values and compute the binary product of these $n$ values.
+This process requires the application of Protocol MYN repeatedly (at least $D$ times) to ensure that every $change$ value is propagated to all $n$ agents.
+Given the importance of this step, which may be performed multiple times during iterations, we first establish a communication framework to propagate the binary product across all $n$ agents effectively.
+    To achieve this, the agents construct a BFS tree of the graph, rooted at the node with the minimum ID agent.
+This BFS tree construction step is a self-independent one-time process and can be completed either just before the first iteration or at the beginning of the decomposition algorithm.
+During the iterations, if the binary product of the $change$ values results in $1$ (indicating that all agents have the same binary product value), the agents terminate the algorithm.
+If not, the agents repeat Phase 3 until the product equals $1$.
+Details of the BFS tree construction mechanism are provided in the subsequent section.
+Creating a BFS Tree for Communication
+As described earlier, to ensure the termination of the Truss Decomposition algorithm, the $n$ agents must communicate to verify that no further edges are scheduled for upgrading and that no additional iterations are required.
+This consensus information must be disseminated across all agents to confirm the completion of the algorithm.
+However, due to the constraints of local communication, propagating such information globally becomes the most time-intensive part of the algorithm.
+To achieve termination through a global consensus, the agents must repeatedly execute Protocol MYN for at least $D$ rounds, where $D$ is the graph’s diameter, each time an edge is flagged for further upgrading.
+The algorithm terminates only after confirming that all agents reach the state $change=1$, indicating that no further computation is needed.
+This process intuitively adds a multiplicative factor of at least $O(D\Delta\log\lambda)$ rounds to the time complexity for each iteration (at-most $m$ times in the worst case), where $m$ is the number of edges.
+The complexity of this process can be significantly reduced if, instead of applying Protocol MYN repeatedly, the agents construct a BFS tree rooted at the minimum ID agent, allowing for more efficient communication across the network.
+In this subsection, we present an algorithm that constructs a BFS tree for the graph $G$, rooted at the agent with the minimum ID, denoted as $r^{\star}$.
+The primary purpose of this BFS tree is to compute a global value $x=f(x_{1},x_{2},x_{3},\dots,x_{n})$ from the individual values $x_{i}$ provided by each agent $r_{i}$, and then disseminate the computed value $x$ to all agents in the network in minimum amounts of time.
+To facilitate the construction of the BFS tree, each agent is equipped with the following variables.
+  • $\bullet$
+    $r_{i}.parent$, $r_{i}.child-$ used for storing respective $child$ and $parent$ pointers for $r_{i}$.
+We recall that each node $u$ with degree $\delta_{u}$ has a local port-numbering =$=\{0,1,\dots,\delta_{u}-1\}$.
+The $r_{i}.parent$ pointer stores the port number that connects $u$ to its parent agent.
+Similarly, the $child$ pointers store the port numbers that connect the node $u$ to its children.
+An agent has at-most $\Delta$ child pointers.
+Initially, $r_{i}.parent,r_{i}.child=\perp$.
+The child pointers uses $O(\Delta\log\Delta)$ bits overall, which is absorbed in the overall memory requirement of $O(\Delta\log n)$ bits for the truss decomposition.
+  • $\bullet$
+    $r_{i}.level-$ denotes the level assigned to an agent relative to the tree that $r_{i}$ is part of. Initially, $r_{i}.level=-1$
+  • $\bullet$
+    $r_{i}.visitor$ a variable set to $1$, if $r_{i}$ is a visitor to another node from its own node. When $r_{i}$ is at its home node, the $visitor$ value is 00.
+The algorithm starts by electing a leader agent, which then coordinates the construction of a BFS tree of the graph $G$.
+After the leader is chosen, the agents proceed to build the BFS tree incrementally, level by level.
+To ensure proper synchronization between levels, the agents utilize the knowledge $\Delta$, the maximum degree of the graph.
+The algorithm operates over $O(D\Delta)$ rounds, with each level of the BFS tree assigned $2\Delta$ rounds for completion.
+Step 1: Electing a Leader
+To elect a leader, each agent $r_{i}$ first sends its ID to its neighbours.
+Each agent now computes the minimum of the IDs received from neighbouring agents and sends this minimum ID to its neighbours.
+Each round of communication with neighbours is performed using Procedure MYN.
+This process continues for $D$ rounds, with each communication requiring $O(\Delta\log\lambda)$ rounds.
+So, to elect a leader an $r_{i}$ does the following steps for $D$ rounds.
+  1. 1.
+    $r_{i}$ calculates the minimum IDs it has received from the neighbours.
+  2. 2.
+    $r_{i}$ sends this minimum ID to its neighbours with one execution of Procedure MYN.
+Using Lemma 4.4, we can argue that after $d$ steps of communication, every agent $r_{i}$, now has the minimum ID of all the agents in its $d-hop$ neighbourhood.
+Since the diameter of the graph is $D$, we have the following lemma.
+Lemma 5.4.
+By the end of $O(D\cdot\Delta\log\lambda)$ rounds every agent has the knowledge of the least ID agent in the network.
+The agents are now aware of the leader agent. We denote this agent as $r^{\star}$. $r^{\star}$ now initiates the construction of the BFS tree rooted at $r^{\star}$.
+Step 2: Construction of BFS Tree
+The leader agent $r^{\star}$ begins constructing a BFS tree, with itself as the root.
+The process proceeds level by level.
+First, $r^{\star}$ marks itself as level 00, initiating the algorithm by visiting all its neighbouring agents one by one.
+As $r^{\star}$ visits each neighbour, it marks them as its children, and these newly marked children also mark $r^{\star}$ as their parent simultaneously.
+To explore the graph in levels, agents use the variable $level$.
+When $r^{\star}$ identifies its children (and they, in turn, recognize $r^{\star}$ as their parent), these agents update their level to $r^{\star}.level+1$.
+The children of $r^{\star}$ then begin exploring the next level in search of their own children.
+However, before they proceed, the children of $r^{\star}$ at level $1$ must ensure that all their siblings have been explored by $r^{\star}$.
+Only after this can agents at the next level begin exploring in parallel.
+This synchronization is crucial to correctly constructing the BFS tree.
+If any child prematurely starts extending the tree before all siblings have been explored, it could result in a tree with longer-than-necessary paths, violating the BFS structure.
+For example, in a complete graph, if the first discovered child immediately continues exploring its own children, the resulting tree could have $n-1$ levels, whereas the correct BFS tree should only have one level.
+To prevent this, the agents synchronize their actions by waiting for $2\Delta$ rounds.
+Below is an outline of the algorithm for an agent $r_{i}$ during rounds $2k\Delta$ to $2(k+1)\Delta$:
+  1. 1.
+    Case 1: $r_{i}.level=k$
+    • $\bullet$
+      $r_{i}$ sets $r_{i}.visitor\leftarrow 1$ moves to its neighbours one by one using increasing port numbers (recall that each node has a local port numbering for every outgoing edge).
+Let one of its neighbours be $v$.
+    • $\bullet$
+      Sub-case 1: If $r_{i}$ meets a single new agent $r_{t}$ at $v$ with $level=-1$ (implying the agent has not been visited before), it marks $r_{t}$ as its child, and simultaneously $r_{t}$ sets $r_{i}$’s incoming port as its parent.
+Additionally, $r_{t}$ sets $r_{t}.level\leftarrow r_{i}.level+1$.
+    • $\bullet$
+      Sub-case 2: $r_{i}$ finds multiple agents at $v$
+      • 1: Original resident present - If the original resident $r_{t}$ of $v$ (with $r_{t}.visitor=0$) is at $v$, $r_{t}$ accepts the request from the visitor agent with the lowest ID.
+If $r_{i}$ is that agent, it updates its child pointers accordingly.
+The newly discovered agent sets its parent accordingly and increments its level to $r_{i}.level+1$.
+If $r_{i}$ is not selected, it returns to its original node and proceeds to the next neighbour.
+      • 2: Original resident absent - If the original resident is absent, all agents are visitors. In this case, the agent $r_{i}$ returns to its original home node.
+    • $\bullet$
+      Sub-case 3: If $r_{i}$ finds $v$ to be empty, $r_{i}$ returns to explore other neighbourhoods.
+An empty node indicates that the resident agent of that particular node is at the same level as $r_{i}$.
+    • $\bullet$
+      Sub-case 4: If $r_{i}$ encounters an agent $r_{t}$ that is already assigned a non-negative level, $r_{i}$ ignores $r_{t}$ and returns to its home node for further exploration.
+    • $\bullet$
+      If all the outgoing ports (neighbours) have been visited, $r_{i}$ sits at its node till the remaining time in $2(k+1)\Delta$ rounds has elapsed.
+  2. 2.
+    Case 2: $r_{i}.level\neq k$: In this case, $r_{i}$ does not take any action and remains idle at its current node during the $2\Delta$ rounds.
+However, if it is visited by a neighbouring agent (which could be a potential parent), $r_{i}$ behaves as described for $r_{t}$ in the previous scenario, responding appropriately to establish the parent-child relationship.
+To complete the BFS tree construction, this process runs in phases of $2\Delta$, a $D$ number of times.
+Therefore, after the completion of $2\Delta D$ rounds, the algorithm terminates and the BFS construction from $r^{\star}$ is complete.
+Since the number of levels in a BFS tree cannot exceed the diameter $D$ of the graph, we have the following lemma.
+Lemma 5.5 (BFS Tree Construction).
+The $n$ agents construct a BFS tree for $G$ rooted at $r^{\star}$ in $O(D\Delta\log\lambda)$ rounds.
+We now provide a mechanism that allows the root $r^{\star}$ to collect $n$ different values $x_{i}$ from each agent $r_{i}$ and compute a value $x=f(x_{1},x_{2},\cdots,x_{n})$ and then disseminate this value to all the agents in $G$.
+  1. 1.
+    Each agent $r_{i}$ sends its own value $x_{i}$ to its parent.
+  2. 2.
+    Each agent waits till it gets the array of values from all of its children, each appends its own value to the array and continues to forward this updated set of data to its parent.
+  3. 3.
+    Finally $r^{\star}$ after receiving all these $x_{i}$ values from its children and using its own value $x^{\star}$ computes the function $x=f(x_{1},x_{2},\cdots,x_{n})$
+  4. 4.
+    To send this value across the BFS tree, the agents use their $level$ number.
+  5. 5.
+    In the first $\Delta$ rounds, the agents with $level=1$ moves to its parent ($r^{\star}$) to collect the value of $x$ from $r^{\star}$.
+  6. 6.
+    In round $t$, each agent $r_{i}$ with $level=t$ similarly moves up to its parent to collect the value of $x$.
+  7. 7.
+    Finally, within $D$ rounds, this value of $x$ reaches every agent in the graph.
+We therefore have the following lemma.
+Lemma 5.6.
+Each agent can learn the value of the function $f(x_{1},x_{2},\dots,x_{n})$, which is computed from $n$ inputs provided by the $n$ agents, in $O(D)$ rounds by utilizing the constructed BFS tree as the framework.
+Therefore, using the methodology above, the agents can compute the binary product of the $change$ values, which is the function $f(x_{1},x_{2},\dots,x_{n})$ here, with $x_{i}$ representing the $change$ values of an agent $r_{i}$.
+In the following lemma, we now analyse the final time complexity of solving the Truss Decomposition Problem using the mobile agents.
+Lemma 5.7.
+Algorithm 5.1 takes $O(m\Delta\log\lambda+mD)$ rounds to terminate.
 Proof.
-Phase 1 and Phase 2 takes $O(\Delta\log\lambda)$ rounds, as shown in Lemma 2 and Lemma 3.
+Phase 1 and Phase 2 take $O(\Delta\log\lambda)$ rounds, as shown in Lemma 4.2 and Lemma 4.3.
+To elect a leader and construct the BFS tree, the agents need additional $O(D\Delta\log\lambda+D\Delta)$ rounds (Lemmas 5.45.5).
 In each execution of Phase 3, the agent needs to communicate with its neighbour to update its $trussness(h)$ values taking $O(\Delta\log\lambda)$ rounds.
-Finally, to check for termination, each agent propagates the $change$ value to every other agent, needing $O(D\Delta\log\lambda)$ rounds.
+Finally, to check for termination, each agent propagates the $change$ value to every other agent, needing $O(D)$ rounds (Lemma 5.6).
 At worst, the algorithm may update the $h-index$ value of one edge during each execution of Phase 3.
@@
 Therefore, the algorithm may need to execute Phase 3, $m$ times before termination.
-Therefore Algorithm 5.1 takes $O(\Delta\log\lambda)+m\cdot(O(\Delta\log\lambda)+O(D\Delta\log\lambda))=O(m\Delta D\log\lambda)$ rounds to execute.
+Therefore Algorithm 5.1 takes $O(\Delta\log\lambda)+O(D\Delta\log\lambda+D\Delta)+m\cdot(O(\Delta\log\lambda)+O(D))=O(m\Delta\log\lambda+mD)$ rounds to execute.
 ∎
@@
 Let $n$ mobile agents with distinct IDs in the range $[0,n^{c}]$ with the highest agent ID $\lambda\in[0,n^{c}]$, where $c$ is constant, be placed at each of the $n$ nodes of $G$ in a dispersed initial configuration.
-Then, the Truss Decomposition Problem for $G$ can be solved by the mobile agents in $O(m\Delta D\log\lambda)$ rounds with $O(\Delta\cdot\log n)$ bits of memory per agent.
+Then, the Truss Decomposition Problem for $G$ can be solved by the mobile agents in $O(m\Delta\log\lambda+mD)$ rounds with $O(\Delta\log n)$ bits of memory per agent.
 5.2 Triangle Centrality
@@
 $TC(v)=\frac{\frac{1}{3}\sum_{u\in N^{+}_{\mathbf{T}}(v)}\mathbf{T}(u)+\sum_{w\in N(v)\setminus N_{\mathbf{T}}(v)}\mathbf{T}(w)}{T(G)}$
-where, where $N(v)$ is the neighborhood set of $v$, $N_{\mathbf{T}}(v)$ is the set of neighbors that are in triangles with $v$, and $N^{+}_{\mathbf{T}}(v)$ is the closed set that includes $v$.
+where, where $N(v)$ is the neighbourhood set of $v$, $N_{\mathbf{T}}(v)$ is the set of neighbours that are in triangles with $v$, and $N^{+}_{\mathbf{T}}(v)$ is the closed set that includes $v$.
 $\mathbf{T}(v)$ and $\mathbf{T}(G)$ denote the respective triangle counts based on $v$ and total triangle count in $G$.
 Mathematically, $N(v)=\{u:(u,v)\in E(G)\}$, $N_{\mathbf{T}}(v)=\{u\in N(v):N(u)\cap N(v)\neq\phi\}$ and $N^{+}_{\mathbf{T}}(v)=\{v\}\cup N_{\mathbf{T}}(v)$.
-Now, we outline our algorithm that computes $TC(v)$ for a node $v\in G$ using the mobile agents that runs in these $3$ phases:
+Now, we outline our algorithm that computes $TC(v)$ for a node $v\in G$ using the mobile agents that run in these $3$ phases:
   1. 1.
     Computing $\mathbf{T}(v)$ and $\mathbf{T}(G)$ We use the triangle counting algorithm for mobile agents described in Section 4 to evaluate the number of triangles involving the node $v$, $\mathbf{T}(v)$ and the total number of triangles in $G$, $\mathbf{T}(G)$.
-The execution of this phase takes $O(\Delta\log\lambda)+O(D\Delta\log\lambda)=O(D\Delta\log\lambda)$ rounds [Lemma 3,6].
+The execution of this phase takes $O(\Delta\log\lambda)+O(D\Delta\log\lambda)=O(D\Delta\log\lambda)$ rounds [Lemma 4.3,4.6].
 With the completion of this phase, the agents begin the next phase.
-Note that if the count of the total number of triangles is available apriori, only $T(v)$ needs to evaluated.
+Note that if the count of the total number of triangles is available a priori, only $T(v)$ needs to be evaluated.
   2. 2.
-    Computing $N(v)$, $N^{+}_{\mathbf{T}}(v)$ and $N_{\mathbf{T}}(v)$ $N(v)$, the neighbours of $v$ can be recorded in $O(\Delta\log\lambda)$ rounds by the agents (Lemma 2).
+    Computing $N(v)$, $N^{+}_{\mathbf{T}}(v)$ and $N_{\mathbf{T}}(v)$ $N(v)$, the neighbours of $v$ can be recorded in $O(\Delta\log\lambda)$ rounds by the agents (Lemma 4.2).
 After we allow the agents to record their neighbourhood in the first $O(\Delta\log\lambda)$ rounds, in the subsequent $O(\Delta\log\lambda)$ rounds, the agents can now meet their neighbours once again, with this time, learning about $N(u)$ for each neighbour agent $u$ of $v$.
@@
 This metric can be easily calculated using mobile agents.
-$T(v)$ for a node $v$ can be calculated in $O(\Delta\log\lambda)$ rounds [Lemma 3].
-The agent which already knows $\delta(v)$ at $v$ can now evaluate $LCC(v)$ using the formula.
+$T(v)$ for a node $v$ can be calculated in $O(\Delta\log\lambda)$ rounds [Lemma 4.3].
+The agent that already knows $\delta(v)$ at $v$ can now evaluate $LCC(v)$ using the formula.
 Theorem 5.10.
@@
 6 Conclusion and Future Work
-In this paper, we introduced algorithms for counting triangles via mobile agents.
-The agents enumerated the triangles based on nodes and edges.
-Finally, the total number of triangles in the graph was calculated.
-We used the triangle enumeration algorithm and applied it to the Truss Decomposition Problem, and to calculate the Triangle Centrality and Local Clustering Coefficient for each node.
-It would interesting to establish lower bounds for the above computations in terms of both time complexity and memory per robot.
-Also, a lower bound on the number of robots used to execute the above algorithm can be an interesting investigation.
-Furthermore, the efficiency of the problem in the faulty robot model can be another task to consider in future.
+In this paper, we introduce and analyse novel algorithms for triangle counting within the mobile agent model over an arbitrary anonymous graph.
+The agents first enumerate triangles based on the nodes and edges, and subsequently calculate the total number of triangles in the graph.
+We subsequently apply this triangle counting methodology to develop and analyse algorithms for solving the Truss Decomposition Problem and for computing important metrics such as Triangle Centrality and Local Clustering Coefficient for each node.
+Future research could focus on several key areas to further enhance our approach.
+Establishing lower bounds for both time complexity and memory usage per agent would provide valuable insights into the efficiency and limitations of our algorithms.
+Additionally, exploring the lower bounds on the number of agents required to execute the algorithms could lead to more optimized and practical implementations.
+Another important direction is to investigate the performance of our algorithms in the presence of faulty agents, which would help in designing more robust solutions for real-world applications.
+Another interesting direction for future work could be to investigate scenarios where global parameters or their bounds, like the maximum degree $\Delta$ or the diameter $D$ of the graph, are not known in advance.